### v1

# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

Abstract
Causal inference typically assumes centralized access to individual-level data. Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints. We address this by estimating the Average Treatment Effect (ATE) from decentralized observational data using federated learning, which enables inference through the exchange of aggregate statistics rather than individual-level data. We propose a novel method to estimate propensity scores in a (non-)parametric manner by computing a federated weighted average of local scores, using two theoretically grounded weighting schemes—Membership Weights (MW) and Density Ratio Weights (DW)—that balance communication efficiency and model flexibility. These federated scores are then used to construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW). Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap. We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions, with theoretical analysis and experiments on simulated and real-world data highlighting their strengths and limitations relative to meta-analysis and related methods.
1 Introduction
In Randomized Clinical Trials (RCTs), treatment assignment is randomized, which ensures that the observed association between treatment and outcome reflects a causal effect. Under this design, the Average Treatment Effect (ATE) can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman et al.,, 1990), which can be further refined through covariate adjustment to reduce variance (U.S. Food and Drug Administration,, 2023; European Medicines Agency,, 2024; Lei and Ding,, 2021). However, RCTs are often expensive, time-consuming, or infeasible in practice. In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán,, 2018; Hernán and Robins,, 2006). Although real-world data is abundant, drawing causal inferences from it remains challenging due to the presence of confounding covariates. As a result, the unadjusted DM estimator is biased when estimating the ATE (Grimes and Schulz,, 2002). To mitigate this bias, it is essential to adjust for confounding variables (VanderWeele,, 2019). This can be done by predicting the counterfactual outcomes of the individuals under the opposite treatment before averaging the differences between the observed and counterfactual outcomes, leading to the so-called G-Formula plug-in estimator (Robins,, 1986). Another popular method for deconfounding is to weight individuals according to their probability of receiving the treatment, thereby emulating a randomized experiment. The Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin,, 1983) relies on the estimation of the propensity score—the probability of treatment given the covariates. Doubly robust estimators, such as the Augmented IPW (AIPW) (Bang and Robins,, 2005), enhance IPW by incorporating an outcome model, offering robustness against misspecification of either model.
In causal inference, larger datasets enhance the precision and reliability of treatment effect estimates, particularly for underrepresented subgroups often overlooked in smaller studies. However, real-world data is typically decentralized—collected across several hospitals, companies, or countries—making it challenging to aggregate into a centralized dataset. This difficulty is particularly acute in domains like healthcare, where privacy concerns, regulatory barriers, data ownership and logistical issues (such as responsibility for data storage and governance) complicate data sharing. Federated Learning (FL) (Kairouz et al.,, 2021) offers a promising solution by enabling model training directly on decentralized data within a server-client architecture, without sharing individual-level data. By exchanging only aggregate statistics such as model updates or gradients, FL has the potential to support collaborative causal inference studies with performance approaching that of centralized analyses. However, realizing this potential requires tailored methodological developments.
Contributions. In this article, we build on ideas from Federated Learning to design (A)IPW estimators for decentralized observational data, moving beyond the simple aggregation of local ATE estimates used in traditional meta-analysis approaches.
At the core of our approach is a flexible, nonparametric strategy for federating propensity scores. Unlike prior methods that aggregate parameters of a global parametric model via “one-shot” averaging (Xiong et al.,, 2023) or federated gradient descent (Khellaf et al.,, 2025), we construct a global propensity score model as a mixture of locally estimated models. This is done through a two-step process: (1) local estimation of propensity scores at each site, which naturally accommodates site-specific heterogeneity in treatment assignment mechanisms and offers flexibility in model choice, and (2) aggregation of these local models into a global one using carefully selected federated weights. We introduce two theoretically grounded weighting schemes: Membership Weights (MW), which represent the probability of an individual belonging to each site given their covariates, and Density Ratio Weights (DW), which model the relative density of covariates at each site compared to the global population. We explain how to infer MW and DW weights in a federated manner and highlight the distinct trade-offs they offer between communication efficiency and modeling assumptions. Using these federated propensity scores, we construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW). We derive asymptotic variance expressions for these estimators and compare them to the meta-analysis estimators—often overlooked in FL (Khellaf et al.,, 2025)––, proving that our estimators achieve lower or equal variances in all scenarios considered.
Our approach is particularly beneficial in scenarios with poor or nonexistent overlap between treated and control groups within individual sites. In such cases, cross-site collaboration becomes crucial, as propensity score functions vary across sites. Indeed, when treatment assignment mechanisms differ substantially—such as when one site treats a subgroup that remains untreated elsewhere—the global dataset achieves markedly greater overlap, providing essential support for estimating treatment effects that would otherwise be poorly estimated within isolated sites. Additionally, our framework naturally accommodates a wide range of cross-site heterogeneities, including variations in sample sizes, treatment policies, covariate distributions, and violations of positivity. Numerical experiments on both simulated and real-world data validate our theoretical findings and highlight the practical value of our method.
2 Preliminaries
2.1 ATE estimators from centralized multi-site observational data
In this section, we recall the key components of ATE estimation in a centralized multi-site setting. Following the potential outcomes framework (Rubin,, 1974; Splawa-Neyman et al.,, 1990), we consider random variables $(X,H,W,Y{(1)},Y{(0)})$, where $X\in\mathbb{R}^{d}$ represents patient covariates, $H\in[K]$ indicates site membership, $W\in\{0,1\}$ denotes the binary treatment, and $Y(1)$ and $Y(0)$ are the potential outcomes under treatment and control, respectively. We assume that the Stable Unit Treatment Values Assumption (SUTVA) holds, so that the observed outcome is $Y=WY(1)+(1-W)Y(0)$, and that potential outcomes are uniformly bounded.
In the centralized setting, we observe $n=\sum_{k=1}^{K}n_{k}$ observations of independently and identically distributed (i.i.d.) tuples $(H_{i},X_{i},W_{i},Y_{i})_{i\in 1,\dots,n}\sim\mathcal{P}$, with $n_{k}=\sum_{i=1}^{n}\mathds{1}_{\{H_{i}=k\}}$ the number of observations in site $k$. We aim to estimate the ATE defined as the risk difference $\tau=\mathbb{E}\left[Y(1)-Y(0)\right]=\mathbb{E}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$, where the expectation is taken over the population $\mathcal{P}$. To be able to identify the ATE, we assume unconfoundedness.
Assumption 1 (Unconfoundedness).
$(Y(0),Y(1))\perp\!\!\!\perp W\mid X$.
While Assumption 1 is a standard assumption in causal inference, we further consider 2, which is specific to the multi-site setting. Together with Assumption 1, it ensures that the variables $X$ form a sufficient set of covariates to adjust for confounding.
Assumption 2 (Ignorability on sites).
$(Y(0),Y(1))\perp\!\!\!\perp H\mid X$.
Our setting is depicted in the graphical model in Figure 1, which highlights that we eliminate any direct effect of the site on the outcome.
$W$$X$$Y$$H$
Figure 1: Graphical model for multi-site observational data.
We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$. The oracle propensity score is denoted by $e(x)=\mathbb{P}(W\mid X=x)$, and we consider the weak (global) overlap assumption (Wager,, 2024).
Assumption 3 (Global overlap).
$\mathcal{O}_{\mathrm{global}}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\big{]}<\infty$.
Assumption 3 is crucial for propensity score-based estimators, as it states that every region of the covariate space has a non-zero probability of receiving both treatments. A lower value of $\mathcal{O}_{\mathrm{global}}$ indicates that these probabilities lie further away from 0 and 1, which corresponds to better overlap. For further insights on overlap, see Li et al., 2018a ; Li et al., 2018b , and for a “misoverlap” metric, refer to Clivio et al., (2024).
With Assumptions 1, 2 and 3, the ATE is identifiable as $\tau=\mathbb{E}\big{[}\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\big{]}$ (see Appendix A.1). We denote oracle ATE estimators, which assume knowledge of the nuisance components $e,\mu_{0},\mu_{1}$, by a superscript . We define the centralized, multi-site estimators as follows.
Definition 1 (Oracle multi-site centralized estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),\quad\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (1)
where $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$ and $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}$.
These oracle estimators are unbiased and asymptotically normal.
Theorem 1 (Consistency of oracle multi-site centralized estimators).
Under Assumptions 1, 2 and 3, we have the following asymptotic normality results:
$\sqrt{n}(\hat{\tau}-\tau)\to\mathcal{N}(0,V^{\star})\ \text{with}$
$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{*}\right]=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},$
$\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{*}\right]=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right],$
where $\tau(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ denotes the Conditional Average Treatment Effect (CATE).
These asymptotic variances align with those in the single-site setting (Hirano et al.,, 2003), as detailed in Appendix A.2. However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data. This creates a challenge in the decentralized setting, where centralizing data to compute $\mu_{1}$, $\mu_{0}$, and $e$ is not feasible. Therefore, the estimators in Definition 1 need to be adapted to this setting. Importantly, the AIPW estimator is doubly robust, meaning it remains consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al.,, 2018).
2.2 Meta-analysis estimators
We now turn to a decentralized setting in which the $K$ sites cannot share individual-level data. A natural baseline for estimating the ATE across sites is a two-stage meta-analysis approach (Burke et al.,, 2017), wherein each site independently estimates the relevant nuisance parameters and communicates only the resulting ATE estimates for aggregation. We will need the following assumption.
Assumption 4 (Local overlap).
$\forall k\in[K],\mathcal{O}_{k}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\mid H=k\big{]}<\infty$.
Assumption 4 is much stronger than its global counterpart (Assumption 3), as it requires the overlap to hold locally for each site. Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows.
Definition 2 (Oracle meta-analysis estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (2)
where $\hat{\tau}_{\mathrm{IPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e_{k})$ and $\hat{\tau}_{\mathrm{AIPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e_{k},\mu_{1},\mu_{0})$ are the local estimators at site $k$.
While alternative aggregation weights—such as the inverse variance of local estimates—can be considered, they produce, in our setting, biased estimates of the global ATE $\tau$, which decomposes as $\tau=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, where $\rho_{k}=\mathbb{P}(H=k)$ and $\tau_{k}=\mathbb{E}\left[Y(1)-Y(0)\mid H=k\right]$ denotes the local ATE at site $k$. This bias arises whenever the $\{\tau_{k}\}_{k}$ differ, which commonly occurs when covariate distributions vary across sites and treatment effects are heterogeneous (i.e., depend on covariates).
Theorem 2 (Consistency of oracle meta-analysis estimators).
Under Assumptions 1, 2 and 4, the oracle meta-analysis estimators are unbiased for the ATE with asymptotic variances
$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right],\quad\quad\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right]$
with within-site variances $V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}$, $V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right]$ and between-sites variance of the local ATEs $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$.
This result is proved in Appendix A.3. A key limitation of meta-analysis estimators is their reliance on Assumption 4, which is fragile and often violated in practice—for example, when a site follows a deterministic treatment policy (e.g., treating all patients or only those with a specific condition). In such cases, these estimators become ill-defined and yield biased estimates of the global ATE.
To overcome this, we introduce a federated approach that estimates the global propensity score $e$ as a weighted combination of the local scores $e_{k}$, enabling valid inference even without local overlap.
3 Related work
Federated causal inference has only recently gained attention. Khellaf et al., (2025) propose a federated framework for estimating the ATE across multiple RCTs by fitting separate parametric linear outcome models for treated and control groups at each site using federated learning, followed by ATE estimation via the G-formula. Their method is motivated by the fact that, in RCTs, this estimator remains unbiased under model misspecification and achieves lower variance than the Difference-in-Means estimator.
In observational studies, however, outcome model misspecification can bias ATE estimates. To address this, Vo et al., (2022) introduce a Bayesian framework using Gaussian processes trained via federated gradient descent. Their method requires estimating a multi-site covariance kernel, which involves each site computing and sharing the first four moments of its data—a computationally intensive process that raises concerns regarding scalability and communication efficiency.
Alternative approaches, such as IPW, require estimating the propensity score. Guo et al., (2025) propose a federated method for learning a global propensity score model using a consensus voting scheme over parametric model parameters, retaining only sites that conform to a shared specification. In contrast, our approach does not assume a common propensity score function: each site can adopt its own, potentially distinct model.
To address site-level heterogeneity, Xiong et al., (2023) assume a parametric logistic propensity score model with shared and site-specific parameters, federating only the common ones in a single round of communication round. In a similar line of work, Yin et al., (2025) estimate a global parametric propensity score model that adjusts for both covariates and site membership, but their handling of heterogeneity is limited to a site-specific scalar, restricting the model’s flexibility. In contrast, our method makes no structural assumptions about the form of the propensity score. We enable fully nonparametric global estimation while allowing flexible and heterogeneous local models, without requiring prior knowledge of shared or site-specific components. Crucially, unlike the above methods, which assume overlap holds locally at each site, our approach explicitly relaxes this strong assumption.
Although not always framed within a federated context, a related body of work focuses on the generalization of causal findings from source data to a target population. Assuming homogeneous outcome models and propensity scores across sites, Han et al., (2025) propose a density ratio weighting method to generalize ATE estimates from multiple source sites to a target domain. Their approach resembles meta-analysis, relying on aggregate statistics—such as sample sizes and local ATE variances—rather than individual-level data. While they estimate density ratios using parametric models, our Membership Weights approach leverages supervised learning methods (e.g., logistic regression) that have lower sample complexity and make less stringent assumptions on the sites’ distributions. Similarly, Guo et al., (2024) address generalization to a target population composed of multiple source sites and previously excluded individuals. Their method assumes the target propensity score is proportional to those of the sources and estimates local scores via density ratios between treated or control subgroups and the target population. However, nonparametric estimation of these ratios requires sufficient data in each treatment arm at every site—a condition often unmet in small or imbalanced settings. Furthermore, the authors do not address how this could be implemented in a federated setting.
4 Federated estimators via propensity score aggregation
4.1 Oracle federated estimators
As discussed in the previous section, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al.,, 2025), site differences limited to intercept shifts (Yin et al.,, 2025), or predefined shared structures (Xiong et al.,, 2023). In practice, treatment assignment often varies across sites due to differences in norms, resources, or clinical practices. To accommodate this heterogeneity, we express the global propensity score as a weighted combination of site-specific scores, capturing local treatment policies.
Specifically, the global propensity score can be expressed as a weighted combination of local scores using two alternative weighting schemes (see Appendix A.4). The first relies on Membership Weights (MW), representing the probability of belonging to each site given the covariates:
$e(X)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X)}_{\text{membership weights}}e_{k}(X).$ (3)
The second uses Density ratio Weights (DW), based on the ratio between the covariate density at site $k$, denoted $f_{k}$, and the overall population density $f$:
$e(X)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(X)}{f(X)}}_{\text{density weights}}e_{k}(X).$ (4)
Building on Eq. 3 and 4, we define our oracle Federated (A)IPW estimators, denoted Fed-(A)IPW, where the corresponding weights will be estimated via federated learning (see Section 4.2).
Definition 3 (Oracle federated estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}},$ (5)
where $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e)$, and $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})$ rely on the global propensity score $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$ with federated weights $\{\omega_{k}(X)\}_{k}$ that are either the membership (Eq. 3) or the density ratio (Eq. 4) weights.
Crucially, our decompositions of the global propensity score in Eq. 3 and 4 enable a flexible combination of locally estimated propensity scores $e_{k}$ and globally learned federation weights $\omega_{k}(x)$. Theorem 3 (proved in Appendix A.5) establishes that this approach, in the oracle setting, matches the efficiency of centralized estimators.
Theorem 3 (Equality of centralized and federated (A)IPW estimators).
Under Assumptions 1, 2 and 3, the oracle federated estimators (Def. 3) are equal to the oracle centralized estimators (Def. 1).
We further show in Appendix A.6 that even when local overlap (Assumption 4) holds, our federated estimators achieve lower variance than meta-analysis estimators.
Theorem 4 (Variance comparison between meta-analysis and federated estimators).
Under Assumptions 1, 2 and 4, both federated and meta-analysis estimators can be computed, and
$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right],$
$\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right],$
with equality when the local propensity scores are equal.
We provide two main reasons for this improvement. First, our decomposition of $e$ in Eq.3 marginalizes over $H\mid X$, effectively removing the need to adjust for site membership $H$. This reduces variance by avoiding unnecessary conditioning. Second, our federated approach inherently yields better overlap than the siloed meta-analysis framework, as formalized in the following theorem.
Theorem 5 (Overlap improvement).
$0\leq\mathcal{O}_{\mathrm{global}}\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}.$
Theorem 5 (proved in Appendix A.7) establishes that global overlap is always at least as good as the worst local overlap. This implies that even when local overlap holds, sites with poor overlap still benefit from the federated approach: the global propensity scores $e(X)$ are more bounded away from 0 and 1 than the local scores $\{e_{k}(X)\}_{k\in[K]}$. Notably, sites with poor individual overlap can even improve the overall overlap of the federated population, as illustrated in the following example.
Example 1. Let $K=2$ with $X_{i}=1$ in both sites, $n_{1}=n_{2}$, $\mathbb{P}(H_{i}\mid X_{i})=0.5$, $e_{1}(X_{i})=0.99\times X_{i}$ and $e_{2}(X_{i})=0.01\times X_{i}$, leading to $e(X_{i})=0.5\times e_{1}(X_{i})+0.5\times e_{2}(X_{i})=0.5\times X_{i}$. Then, $\mathcal{O}_{1}=\mathcal{O}_{2}=\left(0.99\times 0.01\right)^{-1}\approx 101$, while $\mathcal{O}_{\mathrm{global}}=\left(0.5\times 0.5\right)^{-1}=4$. This illustrates how heterogeneity among local propensity scores over a shared population can enhance global overlap.
4.2 Federated estimation
We now move beyond oracle estimators and describe how to implement the Fed-(A)IPW estimators in a practical federated learning setting. Our method follows a two-step procedure (which can be executed in parallel): each site $k$ first estimates and shares a local propensity score function $\hat{e}_{k}(x)$; then, global scores $\hat{e}(x)$ are computed as weighted averages of the local scores, using weights $\{\hat{\omega}_{k}(x)\}_{k\in[K]}$ estimated in a federated manner. For Fed-AIPW, a third step is needed to train the outcome models $\hat{\mu}_{0},\hat{\mu}_{1}$ with federated learning. Details on how to estimate $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$, along with $\hat{\mu}_{0},\hat{\mu}_{1}$, are provided in the following subsections.
4.2.1 Estimation of the local propensity scores
Local propensity scores can be estimated using any probabilistic binary classification method. Common choices include logistic regression for parametric estimation and generalized random forests for non-parametric estimation (Lee et al.,, 2010). A key advantage of our approach is its flexibility: different estimation methods can be used across sites, tailored to local data characteristics or computational constraints.
As previously discussed, this procedure does not require the overlap assumption to hold locally. That is, local propensity scores can legitimately approach 0 or 1—as long as the global score remains sufficiently bounded away from these extremes. This property makes the approach suitable for scenarios involving external control arms (Center for Drug Evaluation and Research and Center for Biologics Evaluation and Research and Oncology Center of Excellence,, 2023; European Medicines Agency,, 2023) where some sites have $e_{k}(X_{i})=0$ for all $i$ in the control group, yet still contribute to the global analysis.
4.2.2 Estimation of the federated weights
Membership weights.
We begin with membership weights $\{\omega_{k}(x)=\mathbb{P}(H_{i}=k\mid X=x)\}_{k}$, which represent the probability of site membership given covariates. These weights can be estimated without access to individual-level data using any federated algorithm for multi-class classification models that can output probabilities.
Federated versions of parametric models—such as logistic regression or deep neural networks—trained via gradient-based empirical risk minimization are well-suited for this task and readily supported in practice (Kairouz et al.,, 2021). Additionally, non-parametric models like random forests (Hauschild et al.,, 2022) and gradient-boosted trees (Li et al.,, 2020) have been adapted to the federated setting, offering more flexibility in capturing complex decision boundaries.
In our experiments, we use a simple multinomial logistic regression model trained via the standard Federated Averaging (FedAvg) algorithm (McMahan et al.,, 2017). This approach is easy to deploy in a server–client architecture, remains interpretable, and requires exchanging only $TKd$ floats, where $T$ is the number of communication rounds. We refer to Appendix B for details.
Density ratios weights.
We now turn to density ratio weights $\{\omega_{k}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}\}_{k}$, where $f_{k}$ and $f$ denote the local and global covariate densities, respectively. Although various methods exist for estimating density ratios (Sugiyama et al.,, 2012), we are not aware of any that are directly compatible with federated settings. Standard non-parametric techniques, such as kernel density estimation, typically require sharing raw data, which violates the decentralized constraints we impose.
We thus propose to estimate each local density $f_{k}$ using a parametric model (e.g. via maximum likelihood estimation). In our experiments, we assume that covariates are normally distributed within each site. Each site $k$ locally estimates its mean $\hat{\mu}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}X_{i}$ and covariance matrix $\hat{\Sigma}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}(X_{i}-\hat{\mu}_{k})(X_{i}-\hat{\mu}_{k})^{\top}$, then transmits these to the central server. Noting that the global density $f$ is a mixture of the $K$ local densities weighted by site proportions $\rho_{k}$ (which can be estimated by $n_{k}/n$), we estimate the density ratios weights as $\hat{\omega}_{k}(x)=\frac{\frac{n_{k}}{n}\hat{f}_{k}(x)}{\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{f}_{k}(x)}$, where each local density is given by
$\hat{f}_{k}(x)$ $=\frac{1}{(2\pi)^{d/2}|\hat{\Sigma}_{k}|^{1/2}}\times\exp\Big{(}-\frac{1}{2}(x-\hat{\mu}_{k})^{\top}\hat{\Sigma}_{k}^{-1}(x-\hat{\mu}_{k})\Big{)}.$
In contrast to the estimation of membership weights, this approach requires only a single round of communication, but involves sending $Kd+Kd^{2}$ floats, which can be significant for high-dimensional data.
More broadly, we recommend using a parametric approach to density ratio estimation when prior knowledge about the local distributions is available and well-justified. However, in the absence of reliable distributional assumptions, membership weights often offer a more robust and practical alternative. They allow for flexible parameterizations and can be estimated through classification models, which are generally easier to learn than densities—particularly in high-dimensional settings where density estimation suffers from high sample complexity and poor scalability (Sugiyama et al.,, 2012).
4.2.3 Estimation of Fed-AIPW
Fed-AIPW requires estimating the outcome models $\mu_{1}$ and $\mu_{0}$ in addition to the propensity components $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$. The estimation procedure for $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$ follows the same steps outlined in Sections 4.2.1 and 4.2.2. For the outcome models, one can estimate $\mu_{0}$ and $\mu_{1}$ globally using federated gradient descent, as in Khellaf et al., (2025), with parametric models such as logistic regression for binary outcomes or linear regression for continuous outcomes. As with the estimation of membership weights, our framework also supports more flexible outcome modeling using a broad range of function classes, including neural networks, decision forests, and other gradient-based models adapted to the federated setting.
5 Experiments
Synthetic data.
We consider $K=3$ sites and $d=10$ covariates. Two distinct data-generating processes (DGPs) are used in the simulations. In DGP A, each site $k$ independently samples $n_{k}=500$ individuals from a site-specific multivariate Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$. In DGP B, a total of $n=1500$ individuals are first drawn from a bimodal Gaussian mixture and then assigned to sites using a multinomial logistic model based on their covariates. We vary the degree of overlap within sites to reflect different practical scenarios: No local overlap (the second site has no treated individuals, hence $\mathcal{O}_{2}=\infty$), Poor local overlap ($\mathcal{O}_{2}\approx 10^{7}$), and Good local overlap ($\mathcal{O}_{2}\approx 4.6$). The outcome models $\mu_{1}$ and $\mu_{0}$ are shared across sites and specified as linear functions with interaction terms. For comparison, we also generate data consistent with the setting described in Xiong et al., (2023) (Figure 5). All reported results are aggregated over 1,500 simulation runs. Complete details of the simulation setup are provided in Appendix C.
We evaluate several estimators: (i) our proposed Fed-IPW and Fed-AIPW, with either Membership Weights (MW) (Def. 3) estimated via federated multinomial logistic regression (well-specified in DGP B), or Density Ratio Weights (DW) (Def. 4) based on Gaussian models (well-specified in DGP A); (ii) the Centralized Oracle (Def. 1); (iii) meta-analysis IPW and AIPW estimators with sample size weighting (Meta-SW) (Def. 2); and (iv) the one-shot inverse variance weighted IPW estimator (1S-IVW) from Xiong et al., (2023), evaluated under a favorable setting for this method—specifically, we assume that certain propensity score parameters are shared across sites. For all estimators, propensity scores are estimated with logistic regression, and outcome models with linear regression.
The results for the three degrees of local overlap are shown in Figures 5, 5 and 5, with the top (respectively bottom) row showing DGP A (respectively DGP B). Before discussing each setting separately, we start by some general observations. First, Fed-IPW-DW is unbiased in DGP A but biased in DGP B, while our Fed-IPW-MW is unbiased in DGP B but biased in DGP A, as expected. Note that Fed-IPW-MW could be made unbiased by using more flexible models (e.g., neural networks) to estimate membership probabilities. Second, our Fed-AIPW estimators remain unbiased across all overlap levels thanks to their double robustness property: despite misspecified linear outcome models, well-estimated propensity scores ensure consistency. Finally, in most cases our Fed-IPW estimators exhibit lower variance than the oracle IPW estimator. This is because they leverage estimated global propensity scores, a strategy known to reduce variance in IPW estimation (Hirano et al.,, 2003).
In the No local overlap setting (Figure 5), meta-analysis estimators are undefined due to the absence of treated individuals at one site. In contrast, our federated estimators remain unbiased under both DGPs, as Assumption 3 holds (with global overlap $\mathcal{O}_{\mathrm{global}}\approx 6.22$). In the Poor local overlap setting (Figure 5), site 2 lacks sufficient overlap, leading to bias and instability in the meta-analysis estimators, including Meta-AIPW as both the propensity weights and local outcome models are inaccurate. However, this issue is mitigated in the global dataset (see Figure 5 in the Appendix), allowing our federated estimators to perform reliably. Finally, in the Good local overlap setting (Figure 5), while all competitors are unbiased, our federated methods exhibit lower variance.
In Figure 5, we consider a specific setting in which all local propensity score functions share a common subset of 5 out of 10 logistic regression coefficients with data generated from DGP B. This setup aligns with the assumptions of the 1S-IVW method proposed by Xiong et al., (2023), which relies on prior knowledge of the shared coefficients—an assumption our method does not require. Nevertheless, Fed-MW is unbiased and exhibits the smallest variance.
Figure 2: Synthetic data: DGP A (top panel) and B (bottom).
(a) No local overlap
(b) Poor local overlap
(c) Good local overlap
Figure 3: Comparison to Xiong et al., (2023)
Figure 4: Results of Traumabase experiment.
Real data.
We apply our methods to the multi-site observational Traumabase cohort, which contains data on patients with traumatic brain injury (Mayer et al.,, 2020; Colnet et al.,, 2024). We study the effect of tranexamic acid (TA), a drug used to reduce bleeding, on the binary outcome of mortality among trauma patients. For this analysis, we select the $K=4$ centres with highest number of treated patients (min. 47), totaling 472 treated and 5,531 control patients, each with 17 measured covariates. We standardize the covariates to have zero mean and unit variance in a federated manner by computing and sharing the means and variances of the covariates across sites.
We focus on AIPW estimators. For Fed-MW and Fed-DW, we estimate the $e_{k}$’s with local logistic regression, while the outcome models $\mu_{1},\mu_{0}$ and the membership weights are learned with a logistic regression model trained with FedAvg (Algorithm 1 in the Appendix) with $T=5000$ rounds, $E=1$ local steps and learning rate $\eta=0.1$. The density ratio weights are computed under a Gaussian model for the covariates. We compare our estimators to Meta-SW and centralized AIPW estimators using also linear and logistic models. To construct empirical confidence intervals, we perform 1000 bootstrap resamples. The results in Figure 4 demonstrate that our federated approach yields estimates comparable to those of the centralized method, while exhibiting lower variance than Meta-SW. Notably, prior analyses using the combined data from a larger set of sites have reported a non-significant average treatment effect (Mayer et al.,, 2020; Colnet et al.,, 2024), which is consistent with our findings on this data subset.
6 Conclusion, limitations and future work
We present a theoretically grounded framework for federated causal inference from decentralized observational data. We introduce two methods for federating local propensity scores—via estimated membership probabilities or density ratios—both of which are designed to be robust to covariate shift and heterogeneity in propensity scores. Our methods improve covariate overlap, leading to more stable and accurate estimates of the ATE. The framework is flexible, allowing the integration of external control arms and supporting heterogeneous local propensity score estimation strategies.
One of the common benefits of federated learning is the ability to increase the overall sample size by leveraging data from multiple sites. While our approach does offer this advantage, it nonetheless requires a sufficiently large amount of data per site, especially when the number of variables is high. Indeed, accurate estimation of key quantities—local propensity scores, parameters of the distribution of the features per site and membership probabilities—relies on sufficient local data. For the latter, the issue becomes more pronounced as the number of sites grows.
RCTs and observational data each have strengths and limitations. Combining them can harness their respective advantages while mitigating weaknesses—supporting goals like improving heterogeneous effect estimation, addressing unobserved confounding, and enhancing generalizability. Estimating treatment effects from diverse sources is thus a promising direction for future federated causal inference. Finally, estimating the CATE in a federated setting is crucial for advancing beyond average treatment effects, towards more personalization in treatment decision-making.
Acknowledgments
This work has been done in the frame of the PEPR SN SMATCH project and has benefited from a governmental grant managed by the Agence Nationale de la Recherche under the France 2030 programme, reference ANR-22-PESN-0003. Rémi Khellaf was supported by an AXIAUM fellowship.
Appendix A Proofs
A.1 ATE Identification and Sufficient Conditions
Proof.
Under Assumptions 1 and 2, the variables in $X_{i}$ are a sufficient adjustment set, so that the ATE is identifiable as
$\tau$ $=\mathbb{E}\left[Y(1)-Y(0)\right]$
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X\right]-\mathbb{E}\left[Y(0)\mid X\right]\right]$
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,W=1\right]-\mathbb{E}\left[Y(0)\mid X,W=0\right]\right]$ Assumptions 1 and 2
$=\mathbb{E}\left[\mathbb{E}\left[Y\mid X,W=1\right]-\mathbb{E}\left[Y\mid X,W=0\right]\right]$ SUTVA
$=\mathbb{E}\left[\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\right]$ Assumption 3
A.2 Proof of Theorem 1
Proof.
Unbiasedness. We have:
$\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\cdot\mathbb{1}_{[H_{i}=k]}\mid n_{1},\dots,n_{K}\right]\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[n_{k}\right]\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$
$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$ (6)
Let us focus on the first term of the sum:
$\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]$ $=\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid H_{i}=k\right]$ SUTVA
$=\mathbb{E}\left[\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid X_{i},H_{i}=k\right]\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[W_{i}\mid X_{i},H_{i}=k\right]\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$ Assumptions 1 and 2
$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$
$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{\mathbb{E}\left[e_{k}(X_{i})\mid X_{i},H_{i}=k\right]}\right]$ Definition of $e(X_{i})$
$=\mathbb{E}\left[\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]\mathbb{E}\left[\frac{e_{k}(X_{i})}{\mathbb{E}\left[e_{k}(X_{i})\right]}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[Y_{i}(1)\mid H_{i}=k\right]$
Similarly, we have $\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]=\mathbb{E}\left[Y_{i}(0)\mid H_{i}=k\right]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
For the AIPW estimator, following the same steps as in section A.2, we have
$\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\sum_{k=1}^{K}\rho_{k}\Bigg{(}\mathbb{E}\left[\mu_{1}(X_{i})-\mu_{0}(X_{i})\mid H_{i}=k\right]$
$\quad\quad+\mathbb{E}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\mid H_{i}=k\right]\Bigg{)}$
$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\tau(X_{i})\mid H_{i}=k\right]+0+0\right)$
$=\tau$
Variance. As we consider uniformly bounded potential outcomes, that is $\forall w\in\{0,1\}$, $\exists(L,U)\in\mathbb{R}^{2},L<Y(w)<U$, and Assumption 3, we have that $\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]<\infty$ and $\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]<\infty$, so the quantities that follow are well defined:
$\mathbb{V}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]$ $=\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{V}\left[\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right]$ i.i.d. observations
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]-\tau^{2}\right)$ unbiasedness
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\right)^{2}\right]+\mathbb{E}\left[\left(\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]-2\mathbb{E}\left[0\right]-\tau^{2}\right)$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\right)^{2}\right]+\mathbb{E}\left[\left(\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]-\tau^{2}\right)$
$\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\right)^{2}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left(\frac{W_{i}Y_{i}}{e(X_{i})}\mid X_{i}\right)^{2}\right]\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[\left(W_{i}Y_{i}\mid X_{i}\right)^{2}\right]}{e(X_{i})^{2}}\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[W_{i}\mid X_{i}\right]\mathbb{E}\left[Y_{i}^{2}\mid X_{i},W_{i}=1\right]}{e(X_{i})^{2}}\right]$ Assumption 1
$=\mathbb{E}\left[\frac{\mathbb{E}\left[Y_{i}(1)^{2}\mid X_{i}\right]}{e(X_{i})}\right]$ SUTVA
$=\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]$
Similarly, $\mathbb{E}\left[\left(\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)^{2}\right]=\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]$. Then, the variance of the oracle multi-site IPW estimator is
$\mathbb{V}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\frac{1}{n}\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]+\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]-\tau^{2}\right).$
For the variance of the oracle multi-site centralized AIPW, we first notice that since $\mathbb{E}\left[Y_{i}(w)-\mu_{w}(X_{i})\mid X_{i}\right]=0$ for treatment $w$, we have
$A$ $=\mathrm{Cov}\left(\tau(X_{i}),\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)$
$=\mathbb{E}\left[\tau(X_{i})\left(\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\mathbb{E}\left[\tau(X_{i})\mathbb{E}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\mid X_{i}\right]\right]-\mathbb{E}\left[\tau(X_{i})\mathbb{E}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[\tau(X_{i})\frac{e(X_{i})\mathbb{E}\left[Y_{i}(1)-\mu_{1}(X_{i})\mid X_{i}\right]}{e(X_{i})}\right]-\mathbb{E}\left[\tau(X_{i})\frac{(1-e(X_{i}))\mathbb{E}\left[Y_{i}(0)-\mu_{0}(X_{i})\mid X_{i}\right]}{1-e(X_{i})}\right]$
$=0$
Then,
$\mathbb{V}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{V}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\right]+\mathbb{V}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right]+2A\right)$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X_{i}))^{2}}{e(X_{i})}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X_{i}))^{2}}{1-e(X_{i})}\right)^{2}\right]\right)$
$=\frac{1}{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X_{i}))^{2}}{e(X_{i})}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X_{i}))^{2}}{1-e(X_{i})}\right)^{2}\right]\right)$
Because these variances are well defined from what precedes, the Central Limit Theorem can be applied to $\hat{\tau}^{\mathrm{IPW^{*}}}$ and $\hat{\tau}^{\mathrm{AIPW^{*}}}$, which gives the result in Theorem 1. ∎
A.3 Proof of Theorem 2
Proof.
Unbiasedness. We first prove the unbiasedness of the local IPW estimators:
$\mathbb{E}\left[\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\mid H_{i}=k\right]$ $=\mathbb{E}\left[\frac{W_{i}Y_{i}}{e_{k}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]$ i.i.d.
$=\mathbb{E}\left[\frac{W_{i}Y_{i}{(1)}}{e_{k}(X_{i})}-\frac{(1-W_{i})Y_{i}{(0)}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]$ SUTVA
$=\mathbb{E}\left[\mathbb{E}\left[\frac{W_{i}Y_{i}{(1)}}{e_{k}(X_{i})}\mid H_{i}=k,X_{i}\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}{(0)}}{1-e_{k}(X_{i})}\mid X_{i},H_{i}=k,X_{i}\right]\right]$
$=\mathbb{E}\Bigg{[}\frac{\mathbb{E}\left[W_{i}\mid H_{i}=k,X_{i}\right]\mathbb{E}\left[Y_{i}{(1)}\mid H_{i}=k,X_{i}\right]}{e_{k}(X_{i})}$
$\quad\quad-\frac{\mathbb{E}\left[(1-W_{i})\mid H_{i}=k,X_{i}\right]\mathbb{E}\left[Y_{i}{(0)}\mid H_{i}=k,X_{i}\right]}{1-e_{k}(X_{i})}\Bigg{]}$ Ass. 1
$=\mathbb{E}\left[\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid H_{i}=k,X_{i}\right]\right]$
$=\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid H_{i}=k\right]$
$=\tau_{k}$
This yields
$\mathbb{E}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ $=\mathbb{E}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\frac{n_{k}}{n}\mathbb{E}\left[\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\mid H_{i}=k\right]\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\frac{n_{k}}{n}\right]\tau_{k}$
$=\sum_{k=1}^{K}\rho_{k}\tau_{k}$
$=\tau$
Variance. For the variance of the local ATEs, we follow the same steps as in the previous proof which yields
$\mathbb{V}\left[\hat{\tau}_{k}^{\mathrm{IPW^{*}}}\mid H=k\right]$ $=\frac{1}{n_{k}}\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau_{k}^{2}\right)$
$=\frac{1}{n_{k}}V_{k,i}$
with $V_{k,i}=\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau_{k}^{2}$. Finally, by Lemma 2, we have
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ $=\mathbb{E}\left[\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\mid H=k\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\mid H=k\right]\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k,i}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right]$
The proof for the meta AIPW estimator follows the same steps. ∎
A.4 Decomposition of the global propensity score
By the law of total probabilities,
$e(x)$ $=\mathbb{P}(W_{i}=1\mid X_{i})$
$=\sum_{k=1}^{K}\mathbb{P}(W_{i}=1\cap H_{i}=k\mid X_{i}=x)$
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})\mathbb{P}(W_{i}=1\mid X_{i}=x,H_{i}=k)$
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})e_{k}(X_{i})$ (Eq. 3)
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\frac{P(X_{i}\mid H_{i}=k)}{P(X_{i})}e_{k}(X_{i})$ (Eq. 4)
A.5 Proof of Theorem 3
Proof.
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}$ $=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}}$
$=\sum_{k=1}^{K}\frac{n_{k}}{n}\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)$
$=\hat{\tau}^{\mathrm{IPW^{*}}}$
We prove similarly that $\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\hat{\tau}^{\mathrm{AIPW^{*}}}$. ∎
A.6 Proof of Theorem 4
We start by two technical lemmas.
Lemma 1.
Under Assumptions 1 and 2 we have,
$\mathbb{V}\left[\tau(X_{i})\right]$ $=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\right]$
Proof.
$\mathbb{V}\left[\tau(X_{i})\right]$ $=\mathbb{E}\left[\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\tau(X_{i})\mid H_{i}=k\right]\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\sum_{k=1}^{K}\mathbb{1}_{[H_{i}=k]}\tau_{H_{i}}\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\sum_{k=1}^{K}\mathbb{1}_{[H_{i}=k]}\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\right]$
Lemma 2.
For the general form of meta-analysis estimator $\hat{\tau}^{\mathrm{meta}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}$, we have
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\frac{1}{n}\mathbb{V}\left[\tau_{H_{i}}\right]$
where $V_{k}=\mathbb{V}\left[\hat{\tau}_{k}\mid H_{i}=k\right]$ is the within-site variance of the ATE estimator in site $k$ and $\mathbb{V}\left[\tau_{H_{i}}\right]=\mathbb{V}\left[\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid H_{i}\right]\right]$ is the between-sites variance of the local ATEs.
Proof.
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta}}\right]$ $=\mathbb{V}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}\right]$
$=\mathbb{E}\left[\mathbb{V}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}\mid H_{1},\dots,H_{n}\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{k}\mid H_{1},\dots,H_{n}\right]\right]$
$=\mathbb{E}\left[\sum_{k=1}^{K}\frac{n_{k}}{n^{2}}V_{k}\right]+\mathbb{V}\left[\sum_{k=1}^{K}\frac{n_{k}}{n}\tau_{k}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\mathbb{V}\left[\frac{1}{n}\sum_{k=1}^{K}\sum_{i=1}^{n}\mathbb{1}_{[H_{i}=k]}\tau_{H_{i}}\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\tau_{H_{i}}\sum_{k=1}^{K}\mathbb{1}_{[H_{i}=k]}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\mathbb{V}\left[\frac{1}{n}\sum_{i=1}^{n}\tau_{H_{i}}\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right],$
We can now prove Theorem 4.
Proof.
We begin with IPW estimators, and then move to AIPW.
IPW.
First, applying Jensen’s inequality with the strictly convex function $t\mapsto\frac{1}{t}$ in ]0;1[]0;1[] 0 ; 1 [ and summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$, we have
$\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]<\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
if $\exists(k,k^{\prime})\in[K],e_{k}(X_{i})\neq e_{k^{\prime}}(X_{i})$, and equality if $\forall k\in[K],e_{k}(X_{i})=e(X_{i})$, i.e. if the local propensity scores are all equal to one another. Then, with the same condition on strictness and equality,
$\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]$ $\leq\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}\mid X_{i}\right]\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid X_{i}\right]\right]$
$=\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
$=\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\mid H_{i}=k\right]$
Similarly, $\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]\leq\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]$. Then,
$\mathbb{V}\left[\hat{\tau}^{\mathrm{fed\text{-}IPW^{*}}}\right]$ $=\frac{1}{n}\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}+\frac{Y_{i}(0)^{2}}{1-e(X_{i})}\right]-\tau^{2}\right)$
$\leq\frac{1}{n}\left(\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}+\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau^{2}\right)$
$\leq\frac{1}{n}\Biggl{(}\sum_{k=1}^{K}\rho_{k}\underbrace{\left(\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}+\frac{Y_{i}(0)^{2}}{1-e_{k}(X_{i})}\mid H_{i}=k\right]-\tau_{k}^{2}\right)}_{:=V_{k}}+\underbrace{\sum_{k=1}^{K}\rho_{k}\tau_{k}^{2}-\tau^{2}}_{\mathbb{V}(\tau_{H})}\Biggr{)}$
On the other hand, by Lemma 2,
$\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{k}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right].$
Hence $\mathbb{V}\left[\hat{\tau}^{\mathrm{fed\text{-}IPW^{*}}}\right]=\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ if $\forall k\in[K],e_{k}=e$, and $\mathbb{V}\left[\hat{\tau}^{\mathrm{fed\text{-}IPW^{*}}}\right]<\mathbb{V}\left[\hat{\tau}^{\mathrm{meta\text{-}IPW^{*}}}\right]$ if $\exists(k,k^{\prime})\in[K]^{2},e_{k}\neq e_{k^{\prime}}$.
AIPW.
Similarly to the IPW case, we have
$\mathbb{E}\left[\left(\frac{W_{i}(Y_{i}-\mu_{1})}{e(X_{i})}\right)^{2}\right]$ $\leq\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\left(\frac{W_{i}(Y_{i}-\mu_{1})}{e_{k}(X_{i})}\right)^{2}\mid H_{i}=k\right],$
$\mathbb{E}\left[\left(\frac{(1-W_{i})(Y_{i}-\mu_{0})}{1-e(X_{i})}\right)^{2}\right]$ $\leq\sum_{k=1}^{K}\rho_{k}\mathbb{E}\left[\left(\frac{(1-W_{i})(Y_{i}-\mu_{0})}{1-e_{k}(X_{i})}\right)^{2}\mid H_{i}=k\right],$
and with Lemma 1:
$\mathbb{V}\left[\tau(X_{i})\right]=\sum_{k=1}^{K}\rho_{k}\mathbb{V}\left[\tau(X_{i})\mid H_{i}=k\right]+\mathbb{V}\left[\tau_{H_{i}}\right].$
Then, using Lemma 2, we have the desired result. ∎
A.7 Proof of Theorem 5
Proof.
Let $f:t\mapsto\frac{1}{t(1-t)}$ with $t\in]0,1[$. $f$ is convex. Then by Jensen’s inequality with summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$,
$\mathcal{O}_{\mathrm{global}}$ $=\mathbb{E}\left[f\left(\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)\right)\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)f(e_{k}(X))\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k\mid X_{i}]}\right]f(e_{k}(X))\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}f(e_{k}(X))\mid X_{i}\right]\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{1}_{[H_{i}=k]}f(e_{k}(X))\right]$
$\leq\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\mathbb{E}\left[f(e_{k}(X))\mid H_{i}=k\right]$
$\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}\qed$
Appendix B Federated learning of membership weights with logistic regression
We detail how to estimate membership weights using a multinomial logistic regression model trained via Federated Averaging (FedAvg) McMahan et al., (2017), without requiring access to individual-level data.
Definition 4 (Multinomial Logistic Model).
The multinomial logistic regression model is defined as
$\omega_{k}(X_{i})=\frac{\exp(\theta_{k}^{\top}X_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\theta_{k^{\prime}}^{\top}X_{i})},$ (7)
where $\theta_{k}$ are the parameters of the model for class $k$.
We introduce the following notations: $\Theta=(\theta_{1},\cdots,\theta_{K})\in\mathbb{R}^{d\times K}$ is the matrix containing the parameters of the multinomial logistic regression, and $\ell(\Theta;\mathcal{D})$ is the negative log-likelihood of the model on a dataset $\mathcal{D}$ given by $\ell(\Theta;\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{k=1}^{K}\mathbb{1}_{\{H_{i}=k\}}\log(\omega_{k}(X_{i}))$ the softmax probabilities. $H_{i}^{\mathrm{enc}}$ is the one-hot encoding of the variable $H_{i}$.
Algorithm 1 presents the FedAvg procedure applied to training a multinomial logistic regression model for estimating membership weights.
Algorithm 1 Federated Multinomial Logistic Regression with FedAvg for Site Membership
1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ rounds of communication, $B$ batch size
2:Initialize model parameters $\Theta_{0}\in\mathbb{R}^{d\times K}$
3:for $t=1$ to $T$ do
4:     for each client $k\in[1,\dots,K]$ in parallel do
5:         $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}^{(k)}$)
6:     end for
7:     $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t}^{(k)}$ // Federated Averaging
8:end for
9:LocalUpdate($k$, $\Theta^{(k)}$, $B$):
10:for $e=1$ to $E$ do
11:     $\mathcal{B}_{k}\leftarrow$ a random batch of $B$ samples from $\mathcal{D}_{k}$
12:     $\omega_{l}(x_{i})\leftarrow\frac{\exp(\theta_{l}^{(k)}{}^{\top}x_{i})}{\sum_{l^{\prime}=1}^{K}\exp(\theta_{l^{\prime}}^{(k)}{}^{\top}x_{i})}$ // Multinomial probabilities
13:     $P_{i}\leftarrow(\omega_{1}(X_{i}),\dots,\omega_{K}(X_{i}))$
14:     $\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})\leftarrow\frac{1}{B}\sum_{i\in\mathcal{B}_{k}}x_{i}(P_{i}-H_{i}^{\mathrm{enc}})$ // Multinomial gradient
15:     $\Theta^{(k)}\leftarrow\Theta^{(k)}-\eta\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})$
16:end for
17:return $\Theta^{(k)}$
With a suitable choice of learning rate $\eta$, and a small number of local steps $E$ per round, Algorithm 1 produces parameter estimates for the multinomial logistic regression model that converge to their centralized counterparts as the number of communication rounds $T$ goes to infinity Stich, (2019); Khaled et al., (2020); Li et al., (2019). After $T$ rounds, Algorithm 1 yields an estimate $\widehat{\Theta}^{\mathrm{fed}}=(\hat{\theta}_{1}^{\mathrm{fed}},\cdots,\hat{\theta}_{K}^{\mathrm{fed}})$ of the multinomial logistic parameters which can be used to estimate the membership weights as:
$\hat{\omega}_{k}(X_{i})=\frac{\exp(\hat{\theta}_{k}^{\mathrm{fed}}{}^{\top}x_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\hat{\theta}_{k^{\prime}}^{\mathrm{fed}}{}^{\top}x_{i})}.$
Appendix C Simulation Details
[…]
Parameter Center 1 Center 2 Center 3
$d$ 10
$\mu_{1}(X)$ $\sum_{j=1}^{5}\frac{j}{10}X_{j}^{2}+\sum_{j=6}^{10}\frac{j}{10}X_{j}+X_{9}*X_{10}$
$\mu_{0}(X)$ $\sum_{j=1}^{5}\frac{3j-10}{30}X_{j}^{2}+\sum_{j=6}^{10}\frac{3j-10}{30}X_{j}+X_{1}*X_{10}$
$e_{k}$ $\mathrm{Logistic}(x,\gamma_{k})$
$\gamma_{k}$ [.$\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$ No overlap: not logistic (only controls) Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$ Good overlap: $\gamma_{2}^{\mathrm{(good)}}$ [.$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
Table 1: Common simulation parameters.
DGP A-specific settings are shown in Table 2, where $J_{d}$ is the $d\times d$ matrix of ones, and $I_{d}$ is the $d\times d$ identity matrix.
Parameter Center 1 Center 2 Center 3
$n_{k}$ 2000
$\mathcal{D}_{k}$ $\mathcal{N}(\mu_{k},\Sigma_{k})$
$\mu_{k}$ $(1,\dots,1)\in\mathbb{R}^{d}$ $(1.5,\dots,1.5)\in\mathbb{R}^{d}$ $(3,\dots,3)\in\mathbb{R}^{d}$
$\Sigma_{k}$ $I_{d}+0.5J_{d}$ $0.6I_{d}+0.4J_{d}$ $3I_{d}+0.3J_{d}$
Table 2: Simulation parameters specific to DGP A.
DGP B-specific settings are shown in Table 3.
Parameter Value
$n$ 6000
$\mathcal{D}$ $\frac{2}{3}\mathcal{N}(\mu_{1},\Sigma_{1})+\frac{1}{3}\mathcal{N}(\mu_{2},\Sigma_{2})$
$\mu_{1}$ $(0,\dots,0)\in\mathbb{R}^{d}$
$\mu_{2}$ $(1.5,\dots,1.5)\in\mathbb{R}^{d}$
$\Sigma_{1}$ $I_{d}$
$\Sigma_{2}$ $I_{d}+0.5J_{d}$
$\mathbb{P}(H_{i}=k\mid X)$ $\mathrm{Logistic}(x,\theta_{k})$
$\theta_{1}$ $[-0.5,-0.5,0.2,-0.5,-0.5,0.2,-0.5,-0.5,0.2,0.2]$
$\theta_{2}$ $[0.5,0.5,0.2,0.5,0.5,0.2,0.5,0.5,0.2,0.5]$
$\theta_{3}$ $[1,1,0.2,0.2,0.2,0.2,0.2,0.2,0.2,0.2]$
Table 3: Simulation parameters specific to DGP B.
Appendix D Overlap Improvement
Echoing theorem 5, Figures 5 and 6 display the empirical distributions (on a log-scale) of the local propensity score in site 2 ($e_{2}$) and of the global propensity score $e$ in the Poor local overlap scenario (see Section 5 for corresponding results). A good overlap is when both the propensity score distributions for the treated and control overlap, and are far from 0. We see that the poor overlap at site 2 (Figure 5), with values of $e_{2}$ on the local data close to 0, is significantly improved at the global level (Figure 6).
Figure 5: Local overlap in site 2 for the Poor local overlap scenario (DGP A).
Figure 6: Global overlap for the Poor local overlap scenario (DGP A). We see a clear improvement compared to the local overlap in site 2 (Figure 5).

### v2

@@
 Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints.
-We address this by estimating the Average Treatment Effect (ATE) from decentralized observational data using federated learning, which enables inference through the exchange of aggregate statistics rather than individual-level data.
-We propose a novel method to estimate propensity scores in a (non-)parametric manner by computing a federated weighted average of local scores, using two theoretically grounded weighting schemes—Membership Weights (MW) and Density Ratio Weights (DW)—that balance communication efficiency and model flexibility.
-These federated scores are then used to construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW).
+We address this problem by estimating the Average Treatment Effect (ATE) from decentralized observational data via a Federated Learning (FL) approach, allowing inference through the exchange of aggregate statistics rather than individual-level data.
+We propose a novel method to estimate propensity scores by computing a federated weighted average of local scores with Membership Weights (MW)—probabilities of site membership conditional on covariates—which can be flexibly estimated using parametric or non-parametric classification models.
+Unlike density ratio weights (DW) from the transportability and generalization literature, which either rely on strong modeling assumptions or cannot be implemented in FL, MW can be estimated using standard FL algorithms and are more robust, as they support flexible, non-parametric models—making them the preferred choice in multi-site settings with strict data-sharing constraints.
+The resulting propensity scores are used to construct Federated Inverse Propensity Weighting (Fed-IPW) and Augmented IPW (Fed-AIPW) estimators.
 Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap.
-We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions, with theoretical analysis and experiments on simulated and real-world data highlighting their strengths and limitations relative to meta-analysis and related methods.
+We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions.
+Both theoretical analysis and experiments on simulated and real-world data highlight their advantages over meta-analysis and related methods.
 1 Introduction
-In Randomized Clinical Trials (RCTs), treatment assignment is randomized, which ensures that the observed association between treatment and outcome reflects a causal effect.
-Under this design, the Average Treatment Effect (ATE) can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman et al.,, 1990), which can be further refined through covariate adjustment to reduce variance (U.S.
-Food and Drug Administration,, 2023; European Medicines Agency,, 2024; Lei and Ding,, 2021).
-However, RCTs are often expensive, time-consuming, or infeasible in practice.
-In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán,, 2018; Hernán and Robins,, 2006).
-Although real-world data is abundant, drawing causal inferences from it remains challenging due to the presence of confounding covariates.
-As a result, the unadjusted DM estimator is biased when estimating the ATE (Grimes and Schulz,, 2002).
-To mitigate this bias, it is essential to adjust for confounding variables (VanderWeele,, 2019).
-This can be done by predicting the counterfactual outcomes of the individuals under the opposite treatment before averaging the differences between the observed and counterfactual outcomes, leading to the so-called G-Formula plug-in estimator (Robins,, 1986).
-Another popular method for deconfounding is to weight individuals according to their probability of receiving the treatment, thereby emulating a randomized experiment.
-The Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin,, 1983) relies on the estimation of the propensity score—the probability of treatment given the covariates.
-Doubly robust estimators, such as the Augmented IPW (AIPW) (Bang and Robins,, 2005), enhance IPW by incorporating an outcome model, offering robustness against misspecification of either model.
-In causal inference, larger datasets enhance the precision and reliability of treatment effect estimates, particularly for underrepresented subgroups often overlooked in smaller studies.
-However, real-world data is typically decentralized—collected across several hospitals, companies, or countries—making it challenging to aggregate into a centralized dataset.
-This difficulty is particularly acute in domains like healthcare, where privacy concerns, regulatory barriers, data ownership and logistical issues (such as responsibility for data storage and governance) complicate data sharing.
-Federated Learning (FL) (Kairouz et al.,, 2021) offers a promising solution by enabling model training directly on decentralized data within a server-client architecture, without sharing individual-level data.
-By exchanging only aggregate statistics such as model updates or gradients, FL has the potential to support collaborative causal inference studies with performance approaching that of centralized analyses.
-However, realizing this potential requires tailored methodological developments.
+The Average Treatment Effect (ATE) is a key causal estimand used to quantify the effect of a treatment on an outcome and is commonly employed as the primary measure of efficacy in evaluating new therapies, including vaccines, before regulatory approval (Polack et al., 2020).
+In Randomized Clinical Trials (RCTs), treatment assignment is randomized, ensuring that the observed association between treatment and outcome reflects a causal effect.
+Under this design, the ATE can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman, 1990), which can be further refined through covariate adjustment to reduce variance (FDA, 2023; EMA, 2024; Lei and Ding, 2021).
+However, RCTs are often expensive, time-consuming, or infeasible.
+In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán, 2018; Hernán and Robins, 2006).
+Although such real-world data is abundant, drawing causal inferences from it is challenging due to confounding covariates, rendering the unadjusted DM estimator biased (Grimes and Schulz, 2002).
+Adjusting for confounders is thus essential (VanderWeele, 2019).
+This can be done by predicting counterfactual outcomes before averaging the differences (the G-formula plug-in estimator, Robins, 1986).
+Another approach is to weight individuals according to their treatment probability, emulating a randomized trial.
+For instance, the Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin, 1983) relies on estimating the propensity score—the probability of treatment given covariates.
+Doubly robust estimators such as the Augmented IPW (AIPW) (Bang and Robins, 2005) combine weighting with outcome modeling to remain consistent as long as either model is correctly specified.
+Larger datasets improve the precision of treatment effect estimates, especially for underrepresented subgroups.
+Yet real-world data is typically decentralized—spread across hospitals, companies, or countries—making aggregation difficult, particularly in healthcare where privacy regulations, data ownership, and governance issues impede centralization.
+Federated Learning (FL) (Kairouz et al., 2021) offers a solution to train models across distributed data without sharing individual-level data.
+While FL has been largely applied to prediction tasks, its extension to causal inference remains limited.
+This problem is especially challenging in observational studies, where differences in covariate distributions and treatment assignment mechanisms across sites create multiple sources of heterogeneity that must be addressed without sharing raw data, while achieving results comparable to centralized analyses—an issue that remains largely unsolved.
 Contributions.
-In this article, we build on ideas from Federated Learning to design (A)IPW estimators for decentralized observational data, moving beyond the simple aggregation of local ATE estimates used in traditional meta-analysis approaches.
-At the core of our approach is a flexible, nonparametric strategy for federating propensity scores.
-Unlike prior methods that aggregate parameters of a global parametric model via “one-shot” averaging (Xiong et al.,, 2023) or federated gradient descent (Khellaf et al.,, 2025), we construct a global propensity score model as a mixture of locally estimated models.
-This is done through a two-step process: (1) local estimation of propensity scores at each site, which naturally accommodates site-specific heterogeneity in treatment assignment mechanisms and offers flexibility in model choice, and (2) aggregation of these local models into a global one using carefully selected federated weights.
-We introduce two theoretically grounded weighting schemes: Membership Weights (MW), which represent the probability of an individual belonging to each site given their covariates, and Density Ratio Weights (DW), which model the relative density of covariates at each site compared to the global population.
-We explain how to infer MW and DW weights in a federated manner and highlight the distinct trade-offs they offer between communication efficiency and modeling assumptions.
-Using these federated propensity scores, we construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW).
-We derive asymptotic variance expressions for these estimators and compare them to the meta-analysis estimators—often overlooked in FL (Khellaf et al.,, 2025)––, proving that our estimators achieve lower or equal variances in all scenarios considered.
-Our approach is particularly beneficial in scenarios with poor or nonexistent overlap between treated and control groups within individual sites.
-In such cases, cross-site collaboration becomes crucial, as propensity score functions vary across sites.
-Indeed, when treatment assignment mechanisms differ substantially—such as when one site treats a subgroup that remains untreated elsewhere—the global dataset achieves markedly greater overlap, providing essential support for estimating treatment effects that would otherwise be poorly estimated within isolated sites.
-Additionally, our framework naturally accommodates a wide range of cross-site heterogeneities, including variations in sample sizes, treatment policies, covariate distributions, and violations of positivity.
-Numerical experiments on both simulated and real-world data validate our theoretical findings and highlight the practical value of our method.
+We propose federated (A)IPW estimators for decentralized observational data, moving beyond the aggregation of local ATE estimates used in meta-analysis (Riley et al., 2023).
+At the core of our approach is a flexible, primarily non-parametric strategy for federating propensity scores.
+Unlike prior methods that fit a single global parametric model via parameter averaging (Xiong et al., 2023) or federated gradient descent (Guo et al., 2025), we construct a global propensity score as a mixture of locally estimated models.
+This involves (1) local estimation of propensity scores at each site—accommodating heterogeneity in treatment assignment and flexibility in model choice—and (2) aggregation into a global score using Membership Weights (MW), i.e., the probability of site membership given covariates.
+MW can be estimated in a federated manner using flexible, potentially non-parametric classification models, ensuring both robustness and communication efficiency.
+In contrast, transferring Density Ratio Weights (DW) from the transportability and generalization literature to our setting requires strong modeling assumptions, as they are otherwise incompatible with FL constraints.
+Using our federated propensity scores, we build the Federated IPW (Fed-IPW) and its augmented variant (Fed-AIPW), derive their variances, and show they achieve equal or lower variance than meta-analysis estimators.
+Our approach is particularly advantageous when overlap between treatment groups is poor or absent within sites.
+In such scenarios, cross-site collaboration becomes crucial, as combining sites increases overall overlap and enables treatment effect estimation that may be infeasible locally.
+Indeed, when treatment assignment mechanisms differ substantially—such as when one site treats a subgroup absent elsewhere—the combined dataset achieves markedly greater overlap, allowing treatment effects to be estimated that would otherwise be poorly identified in isolation.
+Additionally, our framework naturally accommodates heterogeneity in sample sizes, treatment policies, covariate distributions, and violations of positivity.
+Numerical experiments on simulated and real data confirm our theoretical findings and highlight the method’s practical benefits.
+Related work.
+Federated causal inference is still nascent.
+Khellaf et al.
+(2025) estimate federated G-formula ATE estimates across multiple RCTs by fitting parametric outcome models at each site via FL, achieving lower variance than DM.
+In observational settings, Vo et al.
+(2022) propose a federated Bayesian approach using Gaussian processes with a shared covariance kernel, but it requires sharing the first four moments of the data, limiting scalability and efficiency.
+Guo et al.
+(2025) learn a global propensity score via consensus voting over parametric parameters, retaining only sites meeting a shared specification.
+In contrast, our method assumes no common propensity score: each site may fit its own model.
+To address site heterogeneity, Xiong et al.
+(2023) use a logistic propensity model with shared and site-specific parameters, federating only the common ones.
+Yin et al.
+(2025) fit a global model adjusting for covariates and site membership but limit heterogeneity to a site-specific scalar.
+By contrast, our method makes no structural assumptions, enabling fully nonparametric estimation with heterogeneous local models and relaxing the need for local overlap at each site.
+A related body of work focuses on generalizing causal findings from multi-site source populations to a target population.
+Han et al.
+(2025) use density ratio weighting of local ATEs to adjust for covariate shift but assume homogeneous nuisance functions across sites and rely on meta-analysis of aggregate statistics.
+Guo et al.
+(2024) extend this idea by applying density ratio weights to aggregate local propensity scores to construct a target-specific score, which requires density estimation within each treatment arm at each site—demanding large sample sizes per arm for stable estimates.
+In both cases, non-parametric density ratio estimation is infeasible under FL constraints, as it requires sharing raw data or detailed covariate representations (e.g., kernel evaluations or high-dimensional histograms).
+In contrast, our MW-based approach leverages flexible parametric or non-parametric supervised models (e.g., logistic regression, neural networks, gradient-boosted trees) for which federated training is already established, operating without sharing raw data, potentially at lower sample complexity, with full support for heterogeneous nuisance functions, and allowing the natural inclusion of external control arms.
 2 Preliminaries
@@
 In this section, we recall the key components of ATE estimation in a centralized multi-site setting.
-Following the potential outcomes framework (Rubin,, 1974; Splawa-Neyman et al.,, 1990), we consider random variables $(X,H,W,Y{(1)},Y{(0)})$, where $X\in\mathbb{R}^{d}$ represents patient covariates, $H\in[K]$ indicates site membership, $W\in\{0,1\}$ denotes the binary treatment, and $Y(1)$ and $Y(0)$ are the potential outcomes under treatment and control, respectively.
-We assume that the Stable Unit Treatment Values Assumption (SUTVA) holds, so that the observed outcome is $Y=WY(1)+(1-W)Y(0)$, and that potential outcomes are uniformly bounded.
-In the centralized setting, we observe $n=\sum_{k=1}^{K}n_{k}$ observations of independently and identically distributed (i.i.d.) tuples $(H_{i},X_{i},W_{i},Y_{i})_{i\in 1,\dots,n}\sim\mathcal{P}$, with $n_{k}=\sum_{i=1}^{n}\mathds{1}_{\{H_{i}=k\}}$ the number of observations in site $k$.
+Following the potential outcomes framework (Rubin, 1974; Splawa-Neyman, 1990), we consider random variables (X,H,W,Y(1),Y(0))(X,H,W,Y{(1)},Y{(0)}), where $X\in\mathbb{R}^{d}$ represents patient covariates, $H\in[K]$ indicates site membership, $W\in\{0,1\}$ denotes the binary treatment, and $Y(1)$ and $Y(0)$ are the potential outcomes under treatment and control, respectively.
+We assume that the Stable Unit Treatment Values Assumption (SUTVA) holds, so that the observed outcome is Y=WY(1)+(1W)Y(0)Y=WY(1)+(1-W)Y(0), and that the potential outcomes are uniformly bounded.
+In the centralized setting, we observe $n=\sum_{k=1}^{K}n_{k}$ observations of independently and identically distributed (i.i.d.) tuples $(H_{i},X_{i},W_{i},Y_{i})_{i\in 1,\dots,n}\sim\mathcal{P}^{\otimes n}$, with $n_{k}=\sum_{i=1}^{n}\mathds{1}_{\{H_{i}=k\}}$ the number of observations in site $k$.
 We aim to estimate the ATE defined as the risk difference $\tau=\mathbb{E}\left[Y(1)-Y(0)\right]=\mathbb{E}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$, where the expectation is taken over the population $\mathcal{P}$.
-To be able to identify the ATE, we assume unconfoundedness.
+To be able to identify the ATE, we assume unconfoundedness (standard in causal inference) and further consider ˜2, which is specific to the multi-site setting.
 Assumption 1 (Unconfoundedness).
 $(Y(0),Y(1))\perp\!\!\!\perp W\mid X$.
-While Assumption 1 is a standard assumption in causal inference, we further consider 2, which is specific to the multi-site setting.
-Together with Assumption 1, it ensures that the variables $X$ form a sufficient set of covariates to adjust for confounding.
 Assumption 2 (Ignorability on sites).
 $(Y(0),Y(1))\perp\!\!\!\perp H\mid X$.
-Our setting is depicted in the graphical model in Figure 1, which highlights that we eliminate any direct effect of the site on the outcome.
-$W$$X$$Y$$H$
+Robertson et al.
+(2021) refer to ˜2 as the no center-outcome association condition.
+It can be stated as a testable null hypothesis requiring that, for every treatment level $w$ and any pair of centers $(k,k^{\prime})$, $\mathbb{E}[Y\mid X,W=w,H=k]=\mathbb{E}[Y\mid X,W=w,H=k^{\prime}]$.
+Combined with Assumption 1, this ensures that $X$ forms a sufficient set of covariates for confounding adjustment.
+Our setting is depicted in the graphical model in Figure 1, highlighting that we remove any direct effect of the site on the outcome.
+We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$, and let $\tau(x)=\mu_{1}(x)-\mu_{0}(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ be the Conditional Average Treatment Effect (CATE).
+The oracle propensity score is denoted by $e(x)=\mathbb{P}(W\mid X=x)$, and we consider the weak (global) overlap assumption (Wager, 2024).
+Assumption 3 (Global overlap).
+$\mathcal{O}_{\mathrm{global}}=\mathbb{E}\big[(e(X)(1-e(X)))^{-1}\big]<+\infty$.
+WWXXYYHH
 Figure 1: Graphical model for multi-site observational data.
-We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$.
-The oracle propensity score is denoted by $e(x)=\mathbb{P}(W\mid X=x)$, and we consider the weak (global) overlap assumption (Wager,, 2024).
-Assumption 3 (Global overlap).
-$\mathcal{O}_{\mathrm{global}}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\big{]}<\infty$.
-Assumption 3 is crucial for propensity score-based estimators, as it states that every region of the covariate space has a non-zero probability of receiving both treatments.
+˜3 is crucial for propensity score-based estimators, as it states that every region of the covariate space has a non-zero probability of receiving both treatments.
 A lower value of $\mathcal{O}_{\mathrm{global}}$ indicates that these probabilities lie further away from 0 and $1$, which corresponds to better overlap.
-For further insights on overlap, see Li et al., 2018a ; Li et al., 2018b , and for a “misoverlap” metric, refer to Clivio et al., (2024).
+For further insights on overlap, see Li et al.
+(2018a, b), and for a “misoverlap” metric, refer to Clivio et al.
+(2024).
 With Assumptions 1, 2 and 3, the ATE is identifiable as $\tau=\mathbb{E}\big[\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\big]$ (see Appendix A.1).
-We denote oracle ATE estimators, which assume knowledge of the nuisance components $e,\mu_{0},\mu_{1}$, by a superscript .
-We define the centralized, multi-site estimators as follows.
-Definition 1 (Oracle multi-site centralized estimators).
+Throughout the paper, we denote oracle ATE estimators, which assume knowledge of the nuisance components $e,\mu_{0},\mu_{1}$, by a superscript .
+We define the Oracle multi-site centralized estimators as follows:
 $\hat{\tau}_{\mathrm{IPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),\quad\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (1)
@@
 These oracle estimators are unbiased and asymptotically normal.
-Theorem 1 (Consistency of oracle multi-site centralized estimators).
-Under Assumptions 1, 2 and 3, we have the following asymptotic normality results:
-$\sqrt{n}(\hat{\tau}-\tau)\to\mathcal{N}(0,V^{\star})\ \text{with}$
-$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{*}\right]=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},$
-$\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{*}\right]=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right],$
-where $\tau(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ denotes the Conditional Average Treatment Effect (CATE).
-These asymptotic variances align with those in the single-site setting (Hirano et al.,, 2003), as detailed in Appendix A.2.
+Theorem 1.
+Under Assumptions 1, 2 and 3, we have $\sqrt{n}(\hat{\tau}-\tau)\to\mathcal{N}(0,V)$ with
+$\left\{\begin{array}[]{l}V_{\mathrm{IPW}}=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},\\ V_{\mathrm{AIPW}}=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right].\end{array}\right$.
+The above asymptotic variances align with those in the single-site setting (Hirano et al., 2003), as detailed in Appendix A.2.
 However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data.
@@
 Therefore, the estimators in Definition 1 need to be adapted to this setting.
-Importantly, the AIPW estimator is doubly robust, meaning it remains consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al.,, 2018).
+Importantly, the (non-oracle) AIPW estimator is inherently doubly robust, remaining consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al., 2018).
 2.2 Meta-Analysis Estimators
 We now turn to a decentralized setting in which the $K$ sites cannot share individual-level data.
-A natural baseline for estimating the ATE across sites is a two-stage meta-analysis approach (Burke et al.,, 2017), wherein each site independently estimates the relevant nuisance parameters and communicates only the resulting ATE estimates for aggregation.
-We will need the following assumption.
+A natural baseline for estimating the ATE across sites is a two-stage meta-analysis approach (Burke et al., 2017), wherein each site independently estimates the relevant nuisance parameters and communicates only the resulting ATE estimates for aggregation.
+In this setting, we need the following assumption.
 Assumption 4 (Local overlap).
-$\forall k\in[K],\mathcal{O}_{k}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\mid H=k\big{]}<\infty$.
-Assumption 4 is much stronger than its global counterpart (Assumption 3), as it requires the overlap to hold locally for each site.
-Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows.
-Definition 2 (Oracle meta-analysis estimators).
+$\forall k\in[K],\mathcal{O}_{k}=\mathbb{E}\big[(e(X)(1-e(X)))^{-1}\mid H=k\big]<+\infty$.
+Assumption 4 is much stronger than global overlap (˜3), as it must hold at every site.
+Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows:
 $\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (2)
 where $\hat{\tau}_{\mathrm{IPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e_{k})$ and $\hat{\tau}_{\mathrm{AIPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e_{k},\mu_{1},\mu_{0})$ are the local estimators at site $k$.
-While alternative aggregation weights—such as the inverse variance of local estimates—can be considered, they produce, in our setting, biased estimates of the global ATE $\tau$, which decomposes as $\tau=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, where $\rho_{k}=\mathbb{P}(H=k)$ and $\tau_{k}=\mathbb{E}\left[Y(1)-Y(0)\mid H=k\right]$ denotes the local ATE at site $k$.
-This bias arises whenever the $\{\tau_{k}\}_{k}$ differ, which commonly occurs when covariate distributions vary across sites and treatment effects are heterogeneous (i.e., depend on covariates).
-Theorem 2 (Consistency of oracle meta-analysis estimators).
+While alternative aggregation weights—such as the inverse variance of local estimates—can be considered, they produce, in our setting, biased estimates of the global ATE $\tau=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, where $\rho_{k}=\mathbb{P}(H=k)$ and $\tau_{k}=\mathbb{E}[Y(1)-Y(0)\mid H=k]$ is the local ATE.
+This bias appears whenever the $\tau_{k}$} differ, which commonly occurs when covariate distributions vary across sites and treatment effects are heterogeneous (i.e., depend on covariates), see (Berenfeld et al., 2025).
+Theorem 2.
 Under Assumptions 1, 2 and 4, the oracle meta-analysis estimators are unbiased for the ATE with asymptotic variances
-$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right],\quad\quad\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right]$
-with within-site variances $V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}$, $V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right]$ and between-sites variance of the local ATEs $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$.
+$V_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],\quad\quad V_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
+with within-site variance
+$\left\{\begin{array}[]{l}V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}\\ V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right],\end{array}\right$.
+and $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$ the between-sites variance of the local ATEs.
 This result is proved in Appendix A.3.
-A key limitation of meta-analysis estimators is their reliance on Assumption 4, which is fragile and often violated in practice—for example, when a site follows a deterministic treatment policy (e.g., treating all patients or only those with a specific condition).
-In such cases, these estimators become ill-defined and yield biased estimates of the global ATE.
-To overcome this, we introduce a federated approach that estimates the global propensity score $e$ as a weighted combination of the local scores $e_{k}$, enabling valid inference even without local overlap.
-3 Related work
-Federated causal inference has only recently gained attention.
-Khellaf et al., (2025) propose a federated framework for estimating the ATE across multiple RCTs by fitting separate parametric linear outcome models for treated and control groups at each site using federated learning, followed by ATE estimation via the G-formula.
-Their method is motivated by the fact that, in RCTs, this estimator remains unbiased under model misspecification and achieves lower variance than the Difference-in-Means estimator.
-In observational studies, however, outcome model misspecification can bias ATE estimates.
-To address this, Vo et al., (2022) introduce a Bayesian framework using Gaussian processes trained via federated gradient descent.
-Their method requires estimating a multi-site covariance kernel, which involves each site computing and sharing the first four moments of its data—a computationally intensive process that raises concerns regarding scalability and communication efficiency.
-Alternative approaches, such as IPW, require estimating the propensity score.
-Guo et al., (2025) propose a federated method for learning a global propensity score model using a consensus voting scheme over parametric model parameters, retaining only sites that conform to a shared specification.
-In contrast, our approach does not assume a common propensity score function: each site can adopt its own, potentially distinct model.
-To address site-level heterogeneity, Xiong et al., (2023) assume a parametric logistic propensity score model with shared and site-specific parameters, federating only the common ones in a single round of communication round.
-In a similar line of work, Yin et al., (2025) estimate a global parametric propensity score model that adjusts for both covariates and site membership, but their handling of heterogeneity is limited to a site-specific scalar, restricting the model’s flexibility.
-In contrast, our method makes no structural assumptions about the form of the propensity score.
-We enable fully nonparametric global estimation while allowing flexible and heterogeneous local models, without requiring prior knowledge of shared or site-specific components.
-Crucially, unlike the above methods, which assume overlap holds locally at each site, our approach explicitly relaxes this strong assumption.
-Although not always framed within a federated context, a related body of work focuses on the generalization of causal findings from source data to a target population.
-Assuming homogeneous outcome models and propensity scores across sites, Han et al., (2025) propose a density ratio weighting method to generalize ATE estimates from multiple source sites to a target domain.
-Their approach resembles meta-analysis, relying on aggregate statistics—such as sample sizes and local ATE variances—rather than individual-level data.
-While they estimate density ratios using parametric models, our Membership Weights approach leverages supervised learning methods (e.g., logistic regression) that have lower sample complexity and make less stringent assumptions on the sites’ distributions.
-Similarly, Guo et al., (2024) address generalization to a target population composed of multiple source sites and previously excluded individuals.
-Their method assumes the target propensity score is proportional to those of the sources and estimates local scores via density ratios between treated or control subgroups and the target population.
-However, nonparametric estimation of these ratios requires sufficient data in each treatment arm at every site—a condition often unmet in small or imbalanced settings.
-Furthermore, the authors do not address how this could be implemented in a federated setting.
+A key limitation of meta-analysis estimators is their reliance on Assumption 4, which is fragile and often violated—for instance, when a site applies a deterministic treatment policy (treating all patients or only a subgroup).
+In such cases, these estimators are ill-defined, yielding biased ATE estimates.
+To address this, we propose a federated approach that constructs the global propensity score $e$ as a weighted combination of local scores $e_{k}$, enabling valid inference even without local overlap.
 3 Federated Estimators via Propensity Score Aggregation
 3.1 Oracle Federated Estimators
-As discussed in the previous section, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al.,, 2025), site differences limited to intercept shifts (Yin et al.,, 2025), or predefined shared structures (Xiong et al.,, 2023).
-In practice, treatment assignment often varies across sites due to differences in norms, resources, or clinical practices.
-To accommodate this heterogeneity, we express the global propensity score as a weighted combination of site-specific scores, capturing local treatment policies.
-Specifically, the global propensity score can be expressed as a weighted combination of local scores using two alternative weighting schemes (see Appendix A.4).
-The first relies on Membership Weights (MW), representing the probability of belonging to each site given the covariates:
-$e(X)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X)}_{\text{membership weights}}e_{k}(X).$ (3)
-The second uses Density ratio Weights (DW), based on the ratio between the covariate density at site $k$, denoted $f_{k}$, and the overall population density $f$:
-$e(X)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(X)}{f(X)}}_{\text{density weights}}e_{k}(X).$ (4)
-Building on Eq. 3 and 4, we define our oracle Federated (A)IPW estimators, denoted Fed-(A)IPW, where the corresponding weights will be estimated via federated learning (see Section 4.2).
-Definition 3 (Oracle federated estimators).
-$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}},$ (5)
-where $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e)$, and $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})$ rely on the global propensity score $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$ with federated weights $\{\omega_{k}(X)\}_{k}$ that are either the membership (Eq.
-3) or the density ratio (Eq.
-4) weights.
-Crucially, our decompositions of the global propensity score in Eq.
-3 and 4 enable a flexible combination of locally estimated propensity scores $e_{k}$ and globally learned federation weights $\omega_{k}(x)$.
-Theorem 3 (proved in Appendix A.5) establishes that this approach, in the oracle setting, matches the efficiency of centralized estimators.
-Theorem 3 (Equality of centralized and federated (A)IPW estimators).
-Under Assumptions 1, 2 and 3, the oracle federated estimators (Def. 3) are equal to the oracle centralized estimators (Def. 1).
-We further show in Appendix A.6 that even when local overlap (Assumption 4) holds, our federated estimators achieve lower variance than meta-analysis estimators.
-Theorem 4 (Variance comparison between meta-analysis and federated estimators).
-Under Assumptions 1, 2 and 4, both federated and meta-analysis estimators can be computed, and
-$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right],$
-$\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right],$
-with equality when the local propensity scores are equal.
-We provide two main reasons for this improvement.
-First, our decomposition of $e$ in Eq.3 marginalizes over $H\mid X$, effectively removing the need to adjust for site membership $H$.
-This reduces variance by avoiding unnecessary conditioning.
-Second, our federated approach inherently yields better overlap than the siloed meta-analysis framework, as formalized in the following theorem.
+As discussed before, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al., 2025), site differences limited to intercept shifts (Yin et al., 2025), or predefined shared structures (Xiong et al., 2023).
+In practice, treatment assignment frequently varies across sites due to differences in norms, resources, or clinical practices.
+To properly account for this heterogeneity, the global propensity score must be expressed as a weighted combination of the site-specific scores (see Appendix A.4).
+A first choice of weights are the density ratio weights (DW) $\omega_{k}^{\mathrm{DW}}$:
+$e(x)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(x)}{f(x)}}_{=\omega_{k}^{\mathrm{DW}}(x)}e_{k}(x),$ (3)
+where $f_{k}$ and $f$ are the covariate densities locally and globally.
+Similar weights are used in transportability methods that reweight data to match a target population (Han et al., 2023, 2025; Guo et al., 2024), though here the objective is to recover the global propensity score across the super-population defined by the $K$ participating sites.
+Unfortunately, DW estimation in a federated setting requires modeling the $f_{k}$’s, which entails strong distributional assumptions and becomes challenging in high dimensions.
+Instead, we propose the Membership Weights (MW) $\omega_{k}^{\mathrm{MW}}$:
+$e(x)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X=x)}_{\displaystyle=\omega_{k}^{\mathrm{MW}}(x)}e_{k}(x),$ (4)
+which represent the probability of site membership given the covariates.
+Unlike DW, MW do not require explicit density modeling and can be estimated directly via federated parametric (e.g., logistic regression, neural networks) or non-parametric (e.g., gradient-boosted trees) classification models, providing a flexible, communication-efficient alternative and making it the preferred choice for federated settings.
+We refer to Section 3.2 for more details on the federated estimation of MW and its advantages over DW.
+Equations 3 and 4 enable combining locally estimated propensity scores $e_{k}$ into a global propensity score using globally learned weights $\omega_{k}(x)$.
+Building on this decomposition, we define our oracle Federated IPW and AIPW estimators (Fed-(A)IPW).
+We define the Oracle federated estimators as follows:
+$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed}(k)},\qquad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed}(k)},$ (5)
+where $\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed}(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e)$ and $\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed}(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})$ rely on the global propensity score $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$.
+Theorem˜3 (proved in Section˜A.5) establishes that, in the oracle setting, Fed-(A)IPW estimators attain the same efficiency as their centralized counterparts.
+Theorem 3.
+Under Assumptions 1, 2, and 3, the oracle federated estimators (Equation 5) are identical to the oracle centralized estimators (Equation 1).
+Theorem˜4 (proved in Appendix A.6) further shows that even when local overlap (Assumption 4) holds, federated estimators have lower variance than meta-analysis estimators.
+Theorem 4.
+Under Assumptions 1, 2 and 4, we have:
+$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}],\quad\quad\quad\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}],$
+with equality when the local propensity scores are identical across sites.
+This variance reduction arises for two reasons.
+First, decomposing $e$ as a weighted sum of $e_{k}$’s marginalizes over $H\mid X$, eliminating unnecessary adjustment for site membership and thereby reducing variance.
+Second, our federated approach improves overlap compared with meta-analysis, as formalized below.
 Theorem 5 (Overlap improvement).
 $0\leq\mathcal{O}_{\mathrm{global}}\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}.$
-Theorem 5 (proved in Appendix A.7) establishes that global overlap is always at least as good as the worst local overlap.
-This implies that even when local overlap holds, sites with poor overlap still benefit from the federated approach: the global propensity scores $e(X)$ are more bounded away from 0 and 1 than the local scores $\{e_{k}(X)\}_{k\in[K]}$.
+Theorem˜5 (proved in Appendix A.7) shows that global overlap is always at least as good as the worst local overlap.
+Even when local overlap holds, sites with poor overlap benefit from the federated approach because the global score $e(x)$ is more bounded away from 0 and $1$ than the local scores $\{e_{k}(x)\}_{k\in[K]}$.
 Notably, sites with poor individual overlap can even improve the overall overlap of the federated population, as illustrated in the following example.
-Example 1.
-Let $K=2$ with $X_{i}=1$ in both sites, $n_{1}=n_{2}$, $\mathbb{P}(H_{i}\mid X_{i})=0.5$, $e_{1}(X_{i})=0.99\times X_{i}$ and $e_{2}(X_{i})=0.01\times X_{i}$, leading to $e(X_{i})=0.5\times e_{1}(X_{i})+0.5\times e_{2}(X_{i})=0.5\times X_{i}$.
-Then, $\mathcal{O}_{1}=\mathcal{O}_{2}=\left(0.99\times 0.01\right)^{-1}\approx 101$, while $\mathcal{O}_{\mathrm{global}}=\left(0.5\times 0.5\right)^{-1}=4$.
-This illustrates how heterogeneity among local propensity scores over a shared population can enhance global overlap.
+Example.
+Let $K=2$ with $X_{i}=1$ in both sites, $\mathbb{P}(H_{i}\mid X_{i})=0.5$, $e_{1}(X_{i})=0.99$ and $e_{2}(X_{i})=0.01$, leading to $e(X_{i})=\sum_{k=1}^{2}0.5\times e_{k}(X_{i})=0.5$.
+Local overlaps are poor, $\mathcal{O}_{1}=\mathcal{O}_{2}=\left(0.99\times 0.01\right)^{-1}\approx 101$, whereas the global overlap is $\mathcal{O}_{\mathrm{global}}=\left(0.5\times 0.5\right)^{-1}=4$—the optimal value achieved in a randomized trial with $50\%$ treatment probability.
+This illustrates how heterogeneity in treatment assignments can enhance global overlap and enable more robust causal inference.
 3.2 Federated estimation
-We now move beyond oracle estimators and describe how to implement the Fed-(A)IPW estimators in a practical federated learning setting.
-Our method follows a two-step procedure (which can be executed in parallel): each site $k$ first estimates and shares a local propensity score function $\hat{e}_{k}(x)$; then, global scores $\hat{e}(x)$ are computed as weighted averages of the local scores, using weights $\{\hat{\omega}_{k}(x)\}_{k\in[K]}$ estimated in a federated manner.
-For Fed-AIPW, a third step is needed to train the outcome models $\hat{\mu}_{0},\hat{\mu}_{1}$ with federated learning.
-Details on how to estimate $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$, along with $\hat{\mu}_{0},\hat{\mu}_{1}$, are provided in the following subsections.
-4.2.1 Estimation of the local propensity scores
-Local propensity scores can be estimated using any probabilistic binary classification method.
-Common choices include logistic regression for parametric estimation and generalized random forests for non-parametric estimation (Lee et al.,, 2010).
-A key advantage of our approach is its flexibility: different estimation methods can be used across sites, tailored to local data characteristics or computational constraints.
-As previously discussed, this procedure does not require the overlap assumption to hold locally.
-That is, local propensity scores can legitimately approach 0 or 1—as long as the global score remains sufficiently bounded away from these extremes.
-This property makes the approach suitable for scenarios involving external control arms (Center for Drug Evaluation and Research and Center for Biologics Evaluation and Research and Oncology Center of Excellence,, 2023; European Medicines Agency,, 2023) where some sites have $e_{k}(X_{i})=0$ for all $i$ in the control group, yet still contribute to the global analysis.
-4.2.2 Estimation of the federated weights
-Membership weights.
-We begin with membership weights $\{\omega_{k}(x)=\mathbb{P}(H_{i}=k\mid X=x)\}_{k}$, which represent the probability of site membership given covariates.
-These weights can be estimated without access to individual-level data using any federated algorithm for multi-class classification models that can output probabilities.
-Federated versions of parametric models—such as logistic regression or deep neural networks—trained via gradient-based empirical risk minimization are well-suited for this task and readily supported in practice (Kairouz et al.,, 2021).
-Additionally, non-parametric models like random forests (Hauschild et al.,, 2022) and gradient-boosted trees (Li et al.,, 2020) have been adapted to the federated setting, offering more flexibility in capturing complex decision boundaries.
-In our experiments, we use a simple multinomial logistic regression model trained via the standard Federated Averaging (FedAvg) algorithm (McMahan et al.,, 2017).
-This approach is easy to deploy in a server–client architecture, remains interpretable, and requires exchanging only $TKd$ floats, where $T$ is the number of communication rounds.
-We refer to Appendix B for details.
-Density ratios weights.
-We now turn to density ratio weights $\{\omega_{k}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}\}_{k}$, where $f_{k}$ and $f$ denote the local and global covariate densities, respectively.
-Although various methods exist for estimating density ratios (Sugiyama et al.,, 2012), we are not aware of any that are directly compatible with federated settings.
-Standard non-parametric techniques, such as kernel density estimation, typically require sharing raw data, which violates the decentralized constraints we impose.
-We thus propose to estimate each local density $f_{k}$ using a parametric model (e.g.
-via maximum likelihood estimation).
-In our experiments, we assume that covariates are normally distributed within each site.
-Each site $k$ locally estimates its mean $\hat{\mu}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}X_{i}$ and covariance matrix $\hat{\Sigma}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}(X_{i}-\hat{\mu}_{k})(X_{i}-\hat{\mu}_{k})^{\top}$, then transmits these to the central server.
-Noting that the global density $f$ is a mixture of the $K$ local densities weighted by site proportions $\rho_{k}$ (which can be estimated by $n_{k}/n$), we estimate the density ratios weights as $\hat{\omega}_{k}(x)=\frac{\frac{n_{k}}{n}\hat{f}_{k}(x)}{\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{f}_{k}(x)}$, where each local density is given by
-$\hat{f}_{k}(x)$ $=\frac{1}{(2\pi)^{d/2}|\hat{\Sigma}_{k}|^{1/2}}\times\exp\Big{(}-\frac{1}{2}(x-\hat{\mu}_{k})^{\top}\hat{\Sigma}_{k}^{-1}(x-\hat{\mu}_{k})\Big{)}.$
-In contrast to the estimation of membership weights, this approach requires only a single round of communication, but involves sending $Kd+Kd^{2}$ floats, which can be significant for high-dimensional data.
-More broadly, we recommend using a parametric approach to density ratio estimation when prior knowledge about the local distributions is available and well-justified.
-However, in the absence of reliable distributional assumptions, membership weights often offer a more robust and practical alternative.
-They allow for flexible parameterizations and can be estimated through classification models, which are generally easier to learn than densities—particularly in high-dimensional settings where density estimation suffers from high sample complexity and poor scalability (Sugiyama et al.,, 2012).
-4.2.3 Estimation of Fed-AIPW
-Fed-AIPW requires estimating the outcome models $\mu_{1}$ and $\mu_{0}$ in addition to the propensity components $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$.
-The estimation procedure for $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$ follows the same steps outlined in Sections 4.2.1 and 4.2.2.
-For the outcome models, one can estimate $\mu_{0}$ and $\mu_{1}$ globally using federated gradient descent, as in Khellaf et al., (2025), with parametric models such as logistic regression for binary outcomes or linear regression for continuous outcomes.
-As with the estimation of membership weights, our framework also supports more flexible outcome modeling using a broad range of function classes, including neural networks, decision forests, and other gradient-based models adapted to the federated setting.
+We now move beyond oracle estimators and describe how to implement our Fed-(A)IPW estimators in a practical federated learning setting.
+Constructing the global score propensity score requires two steps, which can be executed in parallel: each site $k$ estimates and shares a local propensity score $\hat{e}_{k}(x)$; and the sites collaboratively estimate federated weights $\{\hat{\omega}_{k}(x)\}_{k\in[K]}$.
+Fed-AIPW adds a third step to train outcome models $\hat{\mu}_{0},\hat{\mu}_{1}$ via federated learning.
+We detail how to estimate $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$ and $\hat{\mu}_{0},\hat{\mu}_{1}$ below.
+Local propensity scores.
+Each $e_{k}$ can be estimated using any probabilistic binary classifier, either parametric (e.g., logistic regression or neural networks) or non-parametric ( (e.g., generalized random forests, Lee et al., 2010).
+A key advantage of our approach is flexibility: sites can use different estimation methods tailored to local data or computational constraints.
+It also does not require Assumption 4: local scores may approach 0 or 1 provided the global score remains bounded away from these extremes.This, in particular, enables the integration of external control arms (FDA, 2023; EMA, 2023), where some sites have $e_{k}(X)=0$ for all control patients yet still contribute to the global analysis.
+Federated weights: density ratio vs.
+membership.
+Parametric density ratio weights $\omega_{k}^{\mathrm{DW}}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}$ can be implemented in a one-shot fashion by sharing local density parameters with the server, which then reconstructs the global mixture and computes the weights.
+A common choice is to assume parametric covariate distributions (say, Gaussian), estimate $(\hat{\mu}_{k},\hat{\Sigma}_{k})$ locally, and transmit them once to the server.
+This requires to communicate $O(Kd^{2})$ parameters and is highly sensitive to model misspecification—an issue that becomes critical in high dimensions.
+Nonparametric density estimation would relax these assumptions but is statistically inefficient and does not yet have practical federated implementations.
+In contrast, our membership weights $\omega_{k}^{\mathrm{MW}}(x)=\mathbb{P}(H=k\mid X=x)$ can be learned with any probabilistic multiclass classifier trained federatively—for example, logistic regression for simplicity and interpretability, or neural networks to capture complex nonlinearities.
+Such models are readily supported by modern FL algorithms and software libraries.
+Using the standard FedAvg algorithm (McMahan et al., 2017) requires exchanging $TKP$ floats (training rounds $\times$ sites $\times$ model parameters), which is feasible for models of practical size and a large number of sites.
+Non-parametric classifiers like random forests (Hauschild et al., 2022) and gradient-boosted trees (Li et al., 2020) have also been adapted to the federated setting.
+Outcome models.
+To construct the doubly robust Fed-AIPW estimator, $\mu_{0},\mu_{1}$ are trained federatively, as in Khellaf et al.
+(2025).
+As for MW, standard FL algorithms and librairies can be used to train a wide range of parametric and non-parametric supervised learning models.
+Remark (Missing values).
+Our approach can naturally accommodate missing data.
+Local propensity scores can be estimated consistently with logistic regression or random forests (Jiang et al., 2020; Josse et al., 2024).
+For membership weights, constant imputation combined with federated random forests provides a consistent solution (Le Morvan et al., 2021), whereas density ratio weights would require adapting more complex federated EM algorithms (Dieuleveut et al., 2021; Marfoq et al., 2021).
+Finally, doubly robust estimators with missing data can be obtained via non-parametric federated outcome models (Mayer et al., 2020).
 4 Experiments
 Synthetic data.
-We consider $K=3$ sites and $d=10$ covariates.
-Two distinct data-generating processes (DGPs) are used in the simulations.
-In DGP A, each site $k$ independently samples $n_{k}=500$ individuals from a site-specific multivariate Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$.
-In DGP B, a total of $n=1500$ individuals are first drawn from a bimodal Gaussian mixture and then assigned to sites using a multinomial logistic model based on their covariates.
-We vary the degree of overlap within sites to reflect different practical scenarios: No local overlap (the second site has no treated individuals, hence $\mathcal{O}_{2}=\infty$), Poor local overlap ($\mathcal{O}_{2}\approx 10^{7}$), and Good local overlap ($\mathcal{O}_{2}\approx 4.6$).
-The outcome models $\mu_{1}$ and $\mu_{0}$ are shared across sites and specified as linear functions with interaction terms.
-For comparison, we also generate data consistent with the setting described in Xiong et al., (2023) (Figure 5).
-All reported results are aggregated over 1,500 simulation runs.
-Complete details of the simulation setup are provided in Appendix C.
-We evaluate several estimators: (i) our proposed Fed-IPW and Fed-AIPW, with either Membership Weights (MW) (Def.
-3) estimated via federated multinomial logistic regression (well-specified in DGP B), or Density Ratio Weights (DW) (Def.
-4) based on Gaussian models (well-specified in DGP A); (ii) the Centralized Oracle (Def.
-1); (iii) meta-analysis IPW and AIPW estimators with sample size weighting (Meta-SW) (Def.
-2); and (iv) the one-shot inverse variance weighted IPW estimator (1S-IVW) from Xiong et al., (2023), evaluated under a favorable setting for this method—specifically, we assume that certain propensity score parameters are shared across sites.
-For all estimators, propensity scores are estimated with logistic regression, and outcome models with linear regression.
-The results for the three degrees of local overlap are shown in Figures 5, 5 and 5, with the top (respectively bottom) row showing DGP A (respectively DGP B).
-Before discussing each setting separately, we start by some general observations.
-First, Fed-IPW-DW is unbiased in DGP A but biased in DGP B, while our Fed-IPW-MW is unbiased in DGP B but biased in DGP A, as expected.
-Note that Fed-IPW-MW could be made unbiased by using more flexible models (e.g., neural networks) to estimate membership probabilities.
-Second, our Fed-AIPW estimators remain unbiased across all overlap levels thanks to their double robustness property: despite misspecified linear outcome models, well-estimated propensity scores ensure consistency.
-Finally, in most cases our Fed-IPW estimators exhibit lower variance than the oracle IPW estimator.
-This is because they leverage estimated global propensity scores, a strategy known to reduce variance in IPW estimation (Hirano et al.,, 2003).
-In the No local overlap setting (Figure 5), meta-analysis estimators are undefined due to the absence of treated individuals at one site.
-In contrast, our federated estimators remain unbiased under both DGPs, as Assumption 3 holds (with global overlap $\mathcal{O}_{\mathrm{global}}\approx 6.22$).
-In the Poor local overlap setting (Figure 5), site 2 lacks sufficient overlap, leading to bias and instability in the meta-analysis estimators, including Meta-AIPW as both the propensity weights and local outcome models are inaccurate.
-However, this issue is mitigated in the global dataset (see Figure 5 in the Appendix), allowing our federated estimators to perform reliably.
-Finally, in the Good local overlap setting (Figure 5), while all competitors are unbiased, our federated methods exhibit lower variance.
-In Figure 5, we consider a specific setting in which all local propensity score functions share a common subset of 5 out of 10 logistic regression coefficients with data generated from DGP B.
-This setup aligns with the assumptions of the 1S-IVW method proposed by Xiong et al., (2023), which relies on prior knowledge of the shared coefficients—an assumption our method does not require.
-Nevertheless, Fed-MW is unbiased and exhibits the smallest variance.
-Figure 2: Synthetic data: DGP A (top panel) and B (bottom).
 (a) No local overlap
@@
 (c) Good local overlap
-Figure 3: Comparison to Xiong et al., (2023)
-Figure 4: Results of Traumabase experiment.
+(d) Synthetic data: DGP A (top) and DGP B (bottom).
+We consider $K=3$ sites and $d=10$ covariates.
+Two data-generating processes (DGPs) are used.
+In DGP A, each site $k$ independently samples $n_{k}=650$ individuals from a site-specific multivariate Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$.
+In DGP B, a total of $n=4000$ individuals are first drawn from a bimodal Gaussian mixture and then assigned to sites according to a multinomial logistic model based on their covariates.
+We vary within-site overlap to mimic different practical scenarios: No local overlap ($\mathcal{O}_{2}=+\infty$, the second site has no treated individuals), Poor local overlap ($\mathcal{O}_{2}\approx 10^{7}$), and Good local overlap ($\mathcal{O}_{2}\approx 4.6$).
+The outcome models $\mu_{1},\mu_{0}$ are shared across sites and specified as polynomial functions with interactions.
+For comparison, we also generate data consistent with the setting in Xiong et al.
+(2023) (Figure 2).
+All results are averaged over 800 simulation runs; full details are provided in Appendix C.
+We evaluate our proposed Fed-IPW and Fed-AIPW using the MW weights estimated either via federated multinomial logistic regression (well specified in DGP B) or via a two-layer Neural Network (NN).
+We compare these methods against several competitors: Fed-IPW and Fed-AIPW with the alternative DW weights based on Gaussian density estimation (well specified in DGP A); the Centralized Oracle (Def.
+1); meta-analysis IPW/AIPW with sample-size weighting (Meta-SW) (Def.
+2); and the one-shot inverse-variance weighted IPW estimator (1S-IVW) of Xiong et al.
+(2023), evaluated under favorable conditions with shared propensity-score parameters across sites.
+For all estimators, propensity scores are fit via logistic regression and outcome models via linear regression, implying misspecification in the latter.
+Figures 44 (top: DGP A; bottom: DGP B) summarize the results under the three overlap regimes.
+Before discussing each setting, we highlight several general observations.
+First, Fed-IPW-DW is unbiased only under DGP A, where the Gaussian specification holds, and exhibits bias under DGP B.
+By contrast, Fed-IPW-MW adapts across data-generating processes: with logistic membership weights it is unbiased in DGP B, and with a more flexible neural network classifier it is unbiased for both DGP A and DGP A, reflecting the greater modeling flexibility of MW relative to DW’s restrictive density specification.
+Second, Fed-AIPW enjoys its doubly robust property and remains unbiased across all overlap levels by double robustness; despite misspecified linear outcome models, accurate propensity estimation ensures consistency.
+Finally, Fed-IPW typically attains lower variance than the centralized oracle IPW, consistent with well-documented efficiency gains from using estimated propensity scores (Hirano et al., 2003).
+In the No local overlap setting (Figure 4), meta-analysis estimators are undefined because one site has no treated individuals.
+Our federated estimators remain unbiased under both DGPs, as ˜3 holds (global overlap $\mathcal{O}_{\mathrm{global}}\approx 6.22$).
+In the Poor local overlap setting (Section˜4), site 2 exhibits weak overlap, leading to bias and instability in meta-analysis estimators, including Meta-AIPW, as both the propensity scores and local outcome models are inaccurate.
+This issue is mitigated in the global dataset (see Figure 4 in the Appendix), allowing our federated estimators to remain reliable.
+In the Good local overlap setting (Figure 4), all methods are unbiased, but federated estimators achieve the smallest variance.
+Figure 2 considers a setting where all local propensity scores share a common subset of 5 of 10 logistic regression coefficients with data generated from DGP B.
+This setup matches the assumptions of the 1S-IVW method of Xiong et al.
+(2023), which relies on prior knowledge of the shared parameters to aggregate them—an assumption not required by our method.
+Fed-MW remains unbiased and attains the lowest variance, matching that of 1S-IVW, even without access to the additional information about shared parameters.
+Additional simulations provided in Appendix E examine scenarios with more sites (Table 4), nonparametric estimation (Table 5), and local propensity model misspecification (Table 6), further confirming the robustness of our approach.
+Figure 2: Comparison to Xiong et al. (2023) (IPW).
+Figure 3: AIPW estimates on Traumabase.
 Real data.
-We apply our methods to the multi-site observational Traumabase cohort, which contains data on patients with traumatic brain injury (Mayer et al.,, 2020; Colnet et al.,, 2024).
-We study the effect of tranexamic acid (TA), a drug used to reduce bleeding, on the binary outcome of mortality among trauma patients.
-For this analysis, we select the $K=4$ centres with highest number of treated patients (min.
-47), totaling 472 treated and 5,531 control patients, each with 17 measured covariates.
-We standardize the covariates to have zero mean and unit variance in a federated manner by computing and sharing the means and variances of the covariates across sites.
-We focus on AIPW estimators.
-For Fed-MW and Fed-DW, we estimate the $e_{k}$’s with local logistic regression, while the outcome models $\mu_{1},\mu_{0}$ and the membership weights are learned with a logistic regression model trained with FedAvg (Algorithm 1 in the Appendix) with $T=5000$ rounds, $E=1$ local steps and learning rate $\eta=0.1$.
-The density ratio weights are computed under a Gaussian model for the covariates.
-We compare our estimators to Meta-SW and centralized AIPW estimators using also linear and logistic models.
-To construct empirical confidence intervals, we perform 1000 bootstrap resamples.
-The results in Figure 4 demonstrate that our federated approach yields estimates comparable to those of the centralized method, while exhibiting lower variance than Meta-SW.
-Notably, prior analyses using the combined data from a larger set of sites have reported a non-significant average treatment effect (Mayer et al.,, 2020; Colnet et al.,, 2024), which is consistent with our findings on this data subset.
+We analyze the multi-site Traumabase cohort (Mayer et al., 2020; Colnet et al., 2024) to estimate the effect of tranexamic acid on mortality across $K=14$ centers.
+The local datasets are highly imbalanced in site sizes (106 to 2,092 patients) and treatment arms (e.g., site 11: 4 treated vs.
+121 controls); there are 17 covariates and n=8,248n=8{,}248 patients in total, of whom 638 were treated.
+Covariates are standardized federatively by sharing site means and variances.
+We focus on AIPW estimators: local propensities $e_{k}$ are estimated via logistic regression; outcome models $\mu_{1},\mu_{0}$ are trained with FedAvg logistic regression (5,000 rounds, 1 local epoch, step size $\eta=0.1$); MW are learned with a FedAvg neural network (one hidden layer, 128 units), while DW are based on Gaussian density estimation.
+Competitors include Meta-SW (same local models, computed only on sites with enough treated units—where the number of treated observations exceeds the covariate dimension) and a centralized AIPW benchmark obtained using the R package grf’s probability_forest function (Tibshirani et al., 2018) on the pooled data.
+Empirical confidence intervals are constructed from 150 bootstrap resamples.
+Figure 3 shows that our federated estimators closely match the centralized benchmark and exhibit lower variance than Meta-SW, which in this application departs from the centralized estimate.
+Among federated methods, Fed-MW is closest to the nonparametric centralized estimator—reflecting its more flexible modeling of propensities and membership—whereas Fed-DW is less stable, likely due to noisy per-site covariance estimates for a 17-dimensional Gaussian ($\approx 17^{2}$ parameters) in small sites.
 5 Conclusion, limitations and future work
-We present a theoretically grounded framework for federated causal inference from decentralized observational data.
-We introduce two methods for federating local propensity scores—via estimated membership probabilities or density ratios—both of which are designed to be robust to covariate shift and heterogeneity in propensity scores.
-Our methods improve covariate overlap, leading to more stable and accurate estimates of the ATE.
-The framework is flexible, allowing the integration of external control arms and supporting heterogeneous local propensity score estimation strategies.
-One of the common benefits of federated learning is the ability to increase the overall sample size by leveraging data from multiple sites.
-While our approach does offer this advantage, it nonetheless requires a sufficiently large amount of data per site, especially when the number of variables is high.
-Indeed, accurate estimation of key quantities—local propensity scores, parameters of the distribution of the features per site and membership probabilities—relies on sufficient local data.
-For the latter, the issue becomes more pronounced as the number of sites grows.
-RCTs and observational data each have strengths and limitations.
-Combining them can harness their respective advantages while mitigating weaknesses—supporting goals like improving heterogeneous effect estimation, addressing unobserved confounding, and enhancing generalizability.
-Estimating treatment effects from diverse sources is thus a promising direction for future federated causal inference.
-Finally, estimating the CATE in a federated setting is crucial for advancing beyond average treatment effects, towards more personalization in treatment decision-making.
-Acknowledgments
-This work has been done in the frame of the PEPR SN SMATCH project and has benefited from a governmental grant managed by the Agence Nationale de la Recherche under the France 2030 programme, reference ANR-22-PESN-0003.
-Rémi Khellaf was supported by an AXIAUM fellowship.
+We propose a theoretically grounded framework for federated causal inference that leverages membership weights to construct valid pseudo-populations across silos.
+These weights, estimated via flexible parametric or nonparametric models, improve overlap and yield more stable ATE estimates, without sharing raw data.
+Our framework accommodates heterogeneous local propensity score estimation strategies, supports external control arms, and remains robust to even extreme local treatment–control imbalances.
+Although sufficiently large per-site datasets are still needed—particularly in high-dimensional settings—our approach is especially well suited to a moderate number of large silos, where FL most effectively increases effective sample size.
+Promising avenues for future work include principled handling of site effects (i.e., relaxing Assumption 2) and extending our framework to CATE estimation.
+While the ATE remains central in econometrics, biomedicine, and public policy, considering CATE is important for moving towards personalization.
+Our work lays a foundation for federated CATE estimation: most learners (T-, S-, X-, and R-learners) rely on nuisance quantities such as propensity scores, and our federated estimation framework could be incorporated into these methods, although further work is needed to assess its formal properties and practical performance.
 Appendix A Proofs
@@
 Similarly, we have $\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]=\mathbb{E}\left[Y_{i}(0)\mid H_{i}=k\right]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
-For the AIPW estimator, following the same steps as in section A.2, we have
+For the AIPW estimator, following the same steps as in section˜A.2, we have
 $\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
@@
 $=\frac{1}{n}\left(\mathbb{V}\left[\tau(X_{i})\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X_{i}))^{2}}{e(X_{i})}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X_{i}))^{2}}{1-e(X_{i})}\right)^{2}\right]\right)$
-Because these variances are well defined from what precedes, the Central Limit Theorem can be applied to $\hat{\tau}^{\mathrm{IPW^{*}}}$ and $\hat{\tau}^{\mathrm{AIPW^{*}}}$, which gives the result in Theorem 1.
+Because these variances are well defined from what precedes, the Central Limit Theorem can be applied to $\hat{\tau}^{\mathrm{IPW^{*}}}$ and $\hat{\tau}^{\mathrm{AIPW^{*}}}$, which gives the result in Theorem˜1.
 ∎
@@
 $=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})\mathbb{P}(W_{i}=1\mid X_{i}=x,H_{i}=k)$
-$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})e_{k}(X_{i})$ (Eq. 3)
-$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\frac{P(X_{i}\mid H_{i}=k)}{P(X_{i})}e_{k}(X_{i})$ (Eq. 4)
+$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\frac{P(X_{i}\mid H_{i}=k)}{P(X_{i})}e_{k}(X_{i})$ (DW)
+$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})e_{k}(X_{i})$ (MW)
 A.5 Proof of Theorem 3
@@
 IPW.
-First, applying Jensen’s inequality with the strictly convex function $t\mapsto\frac{1}{t}$ in ]0;1[]0;1[] 0 ; 1 [ and summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$, we have
+First, applying Jensen’s inequality with the strictly convex function $t\mapsto\frac{1}{t}$ in ]0;1[]0;1[ and summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$, we have
 $\mathbb{E}\left[\frac{Y_{i}(1)^{2}}{e(X_{i})}\right]<\mathbb{E}\left[\sum_{k=1}^{K}\omega_{k}(X)\frac{Y_{i}(1)^{2}}{e_{k}(X_{i})}\right]$
@@
 $\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}\qed$
-Appendix B Federated learning of membership weights with logistic regression
-We detail how to estimate membership weights using a multinomial logistic regression model trained via Federated Averaging (FedAvg) McMahan et al., (2017), without requiring access to individual-level data.
-Definition 4 (Multinomial Logistic Model).
-The multinomial logistic regression model is defined as
-$\omega_{k}(X_{i})=\frac{\exp(\theta_{k}^{\top}X_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\theta_{k^{\prime}}^{\top}X_{i})},$ (7)
-where $\theta_{k}$ are the parameters of the model for class $k$.
-We introduce the following notations: $\Theta=(\theta_{1},\cdots,\theta_{K})\in\mathbb{R}^{d\times K}$ is the matrix containing the parameters of the multinomial logistic regression, and $\ell(\Theta;\mathcal{D})$ is the negative log-likelihood of the model on a dataset $\mathcal{D}$ given by $\ell(\Theta;\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{k=1}^{K}\mathbb{1}_{\{H_{i}=k\}}\log(\omega_{k}(X_{i}))$ the softmax probabilities.
-$H_{i}^{\mathrm{enc}}$ is the one-hot encoding of the variable $H_{i}$.
-Algorithm 1 presents the FedAvg procedure applied to training a multinomial logistic regression model for estimating membership weights.
-Algorithm 1 Federated Multinomial Logistic Regression with FedAvg for Site Membership
-1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ rounds of communication, $B$ batch size
-2:Initialize model parameters $\Theta_{0}\in\mathbb{R}^{d\times K}$
+Appendix B Federated learning of membership weights
+We describe how to estimate membership weights using a general parametric classification model trained via Federated Averaging (FedAvg) [McMahan et al., 2017], without requiring access to individual-level data.
+B.1 General Parametric Model
+Let $\omega_{k}(X_{i};\Theta)$ denote the predicted membership probability of site $k$ for covariate vector $X_{i}$, where $\Theta$ are the parameters of the model.
+For a generic parametric model, $\Theta\in\mathbb{R}^{p}$ represents all trainable parameters, and the membership-weight estimation problem can be formulated as the minimization of the empirical cross-entropy loss:
+$\ell(\Theta;\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{k=1}^{K}\mathbb{1}_{\{H_{i}=k\}}\log\bigl(\omega_{k}(X_{i};\Theta)\bigr),$
+where $H_{i}^{\mathrm{enc}}$ is the one-hot encoding of the site membership $H_{i}$.
+This framework encompasses a broad class of models, including:
+  • Multinomial logistic regression:
+    $\omega_{k}(X_{i};\Theta)=\frac{\exp(\theta_{k}^{\top}X_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\theta_{k^{\prime}}^{\top}X_{i})},\quad\Theta=(\theta_{1},\dots,\theta_{K})\in\mathbb{R}^{d\times K}.$
+  • Neural networks, where $\Theta$ includes all weights and biases across layers, and $\omega_{k}(\cdot)$ is the softmax output of the final layer.
+B.2 Federated Training Procedure
+Algorithm 1 summarizes the generic Federated Averaging procedure for training any differentiable parametric model for membership-weight estimation.
+Algorithm 1 Federated Learning of Membership Weights with a Parametric Model
+1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ communication rounds, $B$ batch size
+2:Initialize global parameters $\Theta_{0}$
 3:for $t=1$ to $T$ do
 4:  for each client $k\in[1,\dots,K]$ in parallel do
-5:         $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}^{(k)}$)
+5:   $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}$)
 6:  end for
-7:     $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t}^{(k)}$ // Federated Averaging
+7:  $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t+1}^{(k)}$ // FedAvg aggregation
 8:end for
-9:LocalUpdate($k$, $\Theta^{(k)}$, $B$):
+9:LocalUpdate($k$, $\Theta$):
 10:for $e=1$ to $E$ do
-11:     $\mathcal{B}_{k}\leftarrow$ a random batch of $B$ samples from $\mathcal{D}_{k}$
-12:     $\omega_{l}(x_{i})\leftarrow\frac{\exp(\theta_{l}^{(k)}{}^{\top}x_{i})}{\sum_{l^{\prime}=1}^{K}\exp(\theta_{l^{\prime}}^{(k)}{}^{\top}x_{i})}$ // Multinomial probabilities
-13:     $P_{i}\leftarrow(\omega_{1}(X_{i}),\dots,\omega_{K}(X_{i}))$
-14:     $\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})\leftarrow\frac{1}{B}\sum_{i\in\mathcal{B}_{k}}x_{i}(P_{i}-H_{i}^{\mathrm{enc}})$ // Multinomial gradient
-15:     $\Theta^{(k)}\leftarrow\Theta^{(k)}-\eta\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})$
-16:end for
-17:return $\Theta^{(k)}$
-With a suitable choice of learning rate $\eta$, and a small number of local steps $E$ per round, Algorithm 1 produces parameter estimates for the multinomial logistic regression model that converge to their centralized counterparts as the number of communication rounds $T$ goes to infinity Stich, (2019); Khaled et al., (2020); Li et al., (2019).
-After $T$ rounds, Algorithm 1 yields an estimate $\widehat{\Theta}^{\mathrm{fed}}=(\hat{\theta}_{1}^{\mathrm{fed}},\cdots,\hat{\theta}_{K}^{\mathrm{fed}})$ of the multinomial logistic parameters which can be used to estimate the membership weights as:
-$\hat{\omega}_{k}(X_{i})=\frac{\exp(\hat{\theta}_{k}^{\mathrm{fed}}{}^{\top}x_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\hat{\theta}_{k^{\prime}}^{\mathrm{fed}}{}^{\top}x_{i})}.$
+11:  Sample mini-batch $\mathcal{B}_{k}$ of size $B$ from local data $\mathcal{D}_{k}$
+12:  Compute predictions $\omega_{k}(X_{i};\Theta)$ for $i\in\mathcal{B}_{k}$
+13:  Compute gradient $\nabla\ell(\Theta;\mathcal{B}_{k})$ of cross-entropy loss
+14:  $\Theta\leftarrow\Theta-\eta\nabla\ell(\Theta;\mathcal{B}_{k})$
+15:end for
+16:return $\Theta$
+With a suitable choice of learning rate $\eta$ and a moderate number of local steps $E$ per round, Algorithm 1 converges to the same solution as centralized training as the number of communication rounds $T\to\infty$ [Stich, 2019, Khaled et al., 2020, Li et al., 2019].
+The resulting federated parameter estimate $\widehat{\Theta}^{\mathrm{fed}}$ yields the estimated membership weights:
+$\hat{\omega}_{k}(X_{i})=\omega_{k}(X_{i};\widehat{\Theta}^{\mathrm{fed}}),\qquad k=1,\dots,K.$
+B.3 Example: Multinomial Logistic Regression
+When the model is multinomial logistic regression, $\Theta$ is the matrix of regression coefficients, and the local gradient takes the closed form:
+$\nabla\ell(\Theta;\mathcal{B}_{k})=\frac{1}{B}\sum_{i\in\mathcal{B}_{k}}X_{i}\bigl(\omega(X_{i};\Theta)-H_{i}^{\mathrm{enc}}\bigr),$
+where $\omega(X_{i};\Theta)$ is the vector of softmax probabilities for sample $i$.
 Appendix C Simulation Details
-[…]
+The parameters common to all settings are shown in Table 1, where $\gamma_{2}^{\mathrm{(weak)}}=[-2.5,-1,-0.15,-0.15,0,-0.15,-1,-0.15,-0.15,0]$ and $\gamma_{2}^{\mathrm{(good)}}=[-.05,-.1,-.05,-.1,.05,-.1,-.05,-.1,.05,-.1]$.
 Parameter Center 1 Center 2 Center 3
@@
 $e_{k}$ $\mathrm{Logistic}(x,\gamma_{k})$
-$\gamma_{k}$ [.$\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$ No overlap: not logistic (only controls) Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$ Good overlap: $\gamma_{2}^{\mathrm{(good)}}$ [.$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
+$\gamma_{k}$ $\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$
+No overlap: not logistic
+(only controls)
+Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$
+Good overlap: $\gamma_{2}^{\mathrm{(good)}}$
+$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
 Table 1: Common simulation parameters.
@@
 Parameter Center 1 Center 2 Center 3
-$n_{k}$ 2000
+$n_{k}$ 650
 $\mathcal{D}_{k}$ $\mathcal{N}(\mu_{k},\Sigma_{k})$
-$\mu_{k}$ $(1,\dots,1)\in\mathbb{R}^{d}$ $(1.5,\dots,1.5)\in\mathbb{R}^{d}$ $(3,\dots,3)\in\mathbb{R}^{d}$
+$\mu_{k}$ $(1,\dots,1)\in\mathbb{R}^{d}$ $(1.5,1.5,1.5,1,\dots,1)\in\mathbb{R}^{d}$ $(2,2,2,1,\dots,1)\in\mathbb{R}^{d}$
 $\Sigma_{k}$ $I_{d}+0.5J_{d}$ $0.6I_{d}+0.4J_{d}$ $3I_{d}+0.3J_{d}$
@@
 Parameter Value
-$n$ 6000
+$n$ 4000
 $\mathcal{D}$ $\frac{2}{3}\mathcal{N}(\mu_{1},\Sigma_{1})+\frac{1}{3}\mathcal{N}(\mu_{2},\Sigma_{2})$
@@
 $\mathbb{P}(H_{i}=k\mid X)$ $\mathrm{Logistic}(x,\theta_{k})$
-$\theta_{1}$ $[-0.5,-0.5,0.2,-0.5,-0.5,0.2,-0.5,-0.5,0.2,0.2]$
+$\theta_{1}$ [0.5,0.5,0.2,0.5,0.5,0.2,0.5,0.5,0.2,0.2][-0.5,-0.5,0.2,-0.5,-0.5,0.2,-0.5,-0.5,0.2,0.2]
 $\theta_{2}$ $[0.5,0.5,0.2,0.5,0.5,0.2,0.5,0.5,0.2,0.5]$
@@
 Appendix D Overlap Improvement
-Echoing theorem 5, Figures 5 and 6 display the empirical distributions (on a log-scale) of the local propensity score in site 2 ($e_{2}$) and of the global propensity score $e$ in the Poor local overlap scenario (see Section 5 for corresponding results).
+Echoing theorem˜5, Figures 4 and 5 display the empirical distributions (on a log-scale) of the local propensity score in site 2 ($e_{2}$) and of the global propensity score $e$ in the Poor local overlap scenario (see Section˜4 for corresponding results).
 A good overlap is when both the propensity score distributions for the treated and control overlap, and are far from 0.
@@
 Figure 4: Local overlap in site 2 for the Poor local overlap scenario (DGP A).
-Figure 6: Global overlap for the Poor local overlap scenario (DGP A). We see a clear improvement compared to the local overlap in site 2 (Figure 5).
+Figure 5: Global overlap for the Poor local overlap scenario (DGP A). We see a clear improvement compared to the local overlap in site 2 (Figure˜4).
+Appendix E Additional Simulations
+E.1 Large number of sites
+We further investigate the robustness of our estimators as the number of sites increases.
+Specifically, we simulate $K=20$ sites under weak local overlap (corresponding to the setting of Figure 4 in the main paper) and report the bias and mean squared error (MSE) for both data-generating processes (DGP A and DGP B).
+Table 4: Bias and mean squared error (MSE) under $K=20$ sites with weak local overlap. Bold values indicate unbiasedness or lowest MSE.
+Method Bias (DGP A) MSE (DGP A) Bias (DGP B) MSE (DGP B)
+Oracle IPW 0.00 0.298 0.00 0.130
+Fed. IPW-MW (logistic) 0.48 0.294 0.00 0.095
+Fed. IPW-MW (NN) 0.00 0.240 0.00 0.123
+Fed. IPW-DW 0.00 0.096 0.50 0.450
+Meta-SW IPW -2.09 13.279 -1.77 7.977
+Oracle AIPW 0.00 0.007 0.00 0.049
+Fed. AIPW-MW (logistic) 0.79 0.639 0.00 0.078
+Fed. AIPW-MW (NN) 0.00 0.153 0.00 0.084
+Fed. AIPW-DW 0.00 0.019 0.25 0.194
+Meta-SW AIPW -1.24 1.775 -2.36 18.465
+As shown in Table 4, estimators using neural network membership weights consistently achieve unbiasedness across both DGPs, although with slightly higher variance compared to their well-specified parametric counterparts.
+This is consistent with the behavior of more flexible models, which trade off a small increase in variance for robustness to misspecification.
+When parametric assumptions are correctly specified (e.g., Gaussian density weights in DGP A, logistic membership weights in DGP B), these methods achieve the lowest MSE.
+Overall, the NN-based approach demonstrates robust consistency across DGPs and scales well with the number of sites.
+E.2 Fully model-agnostic ATE estimation
+We additionally investigate the benefits of fully model-agnostic estimation for membership weights $P(H\mid X)$.
+In this setting, we use federated neural networks to estimate membership probabilities in a flexible way and pair them with non-parametric local propensity score estimators (e.g., Random Forests).
+This procedure does not require prior parametric knowledge about $H\mid X$ or $X\mid H$, making it robust to a wide range of DGPs.
+We simulate $K=3$ sites with $d=10$ covariates, $n_{k}=600$ observations per site, and $X\mid H=k\sim D_{k}$ as a mixture of Gaussians with weak local overlap. The results are summarized in Table 5.
+Method Bias MSE
+Oracle IPW 0.00 0.116
+Fed. IPW-MW (NN MW + RF $e_{k}$) 0.00 0.075
+Meta-SW IPW (RF $e_{k}$) 0.89 0.884
+Oracle AIPW 0.00 0.030
+Fed. AIPW-MW (NN MW + RF $e_{k}$) 0.00 0.040
+Meta-SW AIPW (RF $e_{k}$) -0.36 0.180
+Table 5: Bias and MSE under fully non-parametric estimation of membership weights and local propensity scores ($K=3$, weak local overlap).
+Bold values indicate unbiased (or nearly unbiased) estimators.
+As shown in Table 5, our federated approach with neural-network-based membership weights and random-forest local propensities yields nearly unbiased estimates with substantially lower MSE compared to meta-analysis estimators.
+This demonstrates that misspecifications in either the membership weights or the local propensity scores can be effectively mitigated by adopting fully non-parametric methods in a federated learning setting.
+E.3 Robustness to local misspecifications
+Similar to meta-analysis estimators, we argue that inconsistencies in local propensity models can be mitigated as $K$ grows, provided that such inconsistencies are marginal relative to the number of correctly specified sites.
+To illustrate this, we simulate $K$ sites under DGP B, each holding $n_{k}=60$ observations drawn from homogeneous covariate distributions and sharing the same true propensity model (logistic).
+We then intentionally misspecify the propensity model for site $k=1$, setting $\widehat{e}_{1}(x)\equiv 0.1$ for all $x$.
+Table 6 reports the bias and mean squared error (MSE) of several estimators under two scenarios: $K=3$ (few sites) and $K=200$ (many sites).
+Method Bias ($K=3$) MSE ($K=3$) Bias ($K=200$) MSE ($K=200$)
+Oracle IPW 0.00 0.272 0.00 0.002
+Fed. IPW (logistic MW) 0.38 0.344 0.00 0.002
+Meta IPW 2.61 7.902 0.03 0.003
+Oracle AIPW 0.00 0.058 0.00 0.001
+Fed. AIPW (logistic MW) -0.33 0.325 -0.01 0.002
+Meta AIPW -0.51 0.481 -0.07 0.007
+Table 6: Bias and mean squared error (MSE) under local misspecification of site $k=1$ propensity model.
+With only $K=3$ sites, both federated and meta-analysis estimators exhibit substantial bias.
+However, as the number of sites increases, the influence of the single misspecified site becomes negligible, and the bias decreases markedly (e.g., from $0.38$ to $0.00$ for Federated IPW with logistic membership weights).
+As expected, the AIPW estimator retains its double-robustness property even under misspecification of the outcome model, leading to consistently low bias and MSE across settings.

### v3

@@
 We address this problem by estimating the Average Treatment Effect (ATE) from decentralized observational data via a Federated Learning (FL) approach, allowing inference through the exchange of aggregate statistics rather than individual-level data.
-We propose a novel method to estimate propensity scores by computing a federated weighted average of local scores with Membership Weights (MW)—probabilities of site membership conditional on covariates—which can be flexibly estimated using parametric or non-parametric classification models.
-Unlike density ratio weights (DW) from the transportability and generalization literature, which either rely on strong modeling assumptions or cannot be implemented in FL, MW can be estimated using standard FL algorithms and are more robust, as they support flexible, non-parametric models—making them the preferred choice in multi-site settings with strict data-sharing constraints.
+We propose a novel method to estimate propensity scores via a federated weighted average of local scores using Membership Weights (MW), defined as probabilities of site membership conditional on covariates.
+MW can be flexibly estimated with parametric or non-parametric classification models using standard FL algorithms.
 The resulting propensity scores are used to construct Federated Inverse Propensity Weighting (Fed-IPW) and Augmented IPW (Fed-AIPW) estimators.
-Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap.
+In contrast to meta-analysis methods, which fail when any site violates positivity, our approach exploits heterogeneity in treatment assignment across sites to improve overlap.
 We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions.
-Both theoretical analysis and experiments on simulated and real-world data highlight their advantages over meta-analysis and related methods.
+Theoretical analysis and experiments on simulated and real-world data demonstrate clear advantages over meta-analysis and related approaches.
 1 Introduction
@@
 In Randomized Clinical Trials (RCTs), treatment assignment is randomized, ensuring that the observed association between treatment and outcome reflects a causal effect.
-Under this design, the ATE can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman, 1990), which can be further refined through covariate adjustment to reduce variance (FDA, 2023; EMA, 2024; Lei and Ding, 2021).
+Under this design, the ATE can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman, 1990), which can be further refined through covariate adjustment to reduce variance (FDA, 2023a; EMA, 2024; Lei and Ding, 2021).
 However, RCTs are often expensive, time-consuming, or infeasible.
@@
 By contrast, our method makes no structural assumptions, enabling fully nonparametric estimation with heterogeneous local models and relaxing the need for local overlap at each site.
-A related body of work focuses on generalizing causal findings from multi-site source populations to a target population.
+A related line of work focuses on generalizing causal findings from multi-site source populations to a target population.
 Han et al.
@@
 Guo et al.
-(2024) extend this idea by applying density ratio weights to aggregate local propensity scores to construct a target-specific score, which requires density estimation within each treatment arm at each site—demanding large sample sizes per arm for stable estimates.
-In both cases, non-parametric density ratio estimation is infeasible under FL constraints, as it requires sharing raw data or detailed covariate representations (e.g., kernel evaluations or high-dimensional histograms).
-In contrast, our MW-based approach leverages flexible parametric or non-parametric supervised models (e.g., logistic regression, neural networks, gradient-boosted trees) for which federated training is already established, operating without sharing raw data, potentially at lower sample complexity, with full support for heterogeneous nuisance functions, and allowing the natural inclusion of external control arms.
+(2024) extend this approach by applying density-ratio weights to combine local propensity scores into a target-specific score.
+While their framework can target the super-population of the $K$ sites as a special case, it is significantly more constrained than our approach: it requires access to individual-level target covariates via an external public dataset, reweights each treatment arm separately within each site (needing large per-arm sample sizes), and relies on direct estimation of density ratios either through strongly parametric logistic models—which are difficult to validate and sensitive to misspecification—or via nonparametric approaches (e.g., kernel-based methods) that involve sharing raw covariates, making them incompatible with federated learning.
+To summarize, our approach stands out for its flexibility and ease of implementation in a federated context: (i) It allows local propensity scores to take arbitrary forms and vary freely across sites, unlike Xiong et al.
+(2023); Han et al.
+(2025), which rely on strong parametric assumptions, sometimes with shared parameters; (ii) It truly satisfies the federated learning requirement that no raw data be shared, unlike density ratio-based methods that are either highly parametric or lack federated algorithms (Yin et al., 2025; Han et al., 2025; Guo et al., 2024); and (iii) It does not require local overlap, unlike Xiong et al.
+(2023); Han et al.
+(2025), and to the best of our knowledge, is the first to provably improve overlap relative to meta-analysis.
 2 Preliminaries
@@
 Assumption 1 (Unconfoundedness).
-$(Y(0),Y(1))\perp\!\!\!\perp W\mid X$.
+$\{Y(0),Y(1)\}\perp\!\!\!\perp W\mid(X,H).$
 Assumption 2 (Ignorability on sites).
-$(Y(0),Y(1))\perp\!\!\!\perp H\mid X$.
-Robertson et al.
-(2021) refer to ˜2 as the no center-outcome association condition.
-It can be stated as a testable null hypothesis requiring that, for every treatment level $w$ and any pair of centers $(k,k^{\prime})$, $\mathbb{E}[Y\mid X,W=w,H=k]=\mathbb{E}[Y\mid X,W=w,H=k^{\prime}]$.
-Combined with Assumption 1, this ensures that $X$ forms a sufficient set of covariates for confounding adjustment.
-Our setting is depicted in the graphical model in Figure 1, highlighting that we remove any direct effect of the site on the outcome.
+$\{Y(0),Y(1)\}\perp\!\!\!\perp H\mid X.$
+We provide a federated test for ˜2 in Section˜F.1.
+With ˜1, they imply $\{Y(0),Y(1)\}\perp H\mid(X,W)$ (Dawid, 1979, Lemma 4.3).
+This entails the no center-outcome association condition (Robertson et al., 2021), yielding the testable null hypothesis $\mathbb{E}[Y\mid X,W,H]=\mathbb{E}[Y\mid X,W]$.
+This ensures that $X$ forms a sufficient set of covariates for confounding adjustment.
+Our setting is depicted in the graphical model in Figure 1, highlighting that we exclude any direct effect of the site on the outcome.
+In Section˜3.3, we introduce a relaxation of ˜2 and extend our approach to handle the presence of site effects.
 We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$, and let $\tau(x)=\mu_{1}(x)-\mu_{0}(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ be the Conditional Average Treatment Effect (CATE).
@@
 We define the Oracle multi-site centralized estimators as follows:
-$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),\quad\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (1)
-where $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$ and $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}$.
-These oracle estimators are unbiased and asymptotically normal.
+$\hat{\tau}_{\mathrm{IPW}}^{*}$ $=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),$ (1)
+$\hat{\tau}_{\mathrm{AIPW}}^{*}$ $=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (2)
+where $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}\!\left(Y_{i}-\mu_{1}(X_{i})\right)}{e(X_{i})}-\frac{(1-W_{i})\!\left(Y_{i}-\mu_{0}(X_{i})\right)}{1-e(X_{i})}$ and $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$.
+These estimators are unbiased and asymptotically normal.
 Theorem 1.
@@
 $\left\{\begin{array}[]{l}V_{\mathrm{IPW}}=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},\\ V_{\mathrm{AIPW}}=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right].\end{array}\right$.
-The above asymptotic variances align with those in the single-site setting (Hirano et al., 2003), as detailed in Appendix A.2.
+The above asymptotic variances align with the single-site setting (Hirano et al., 2003), as detailed in Appendix A.2.
 However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data.
@@
 Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows:
-$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (2)
+$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (3)
 where $\hat{\tau}_{\mathrm{IPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e_{k})$ and $\hat{\tau}_{\mathrm{AIPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e_{k},\mu_{1},\mu_{0})$ are the local estimators at site $k$.
@@
 Under Assumptions 1, 2 and 4, the oracle meta-analysis estimators are unbiased for the ATE with asymptotic variances
-$V_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],\quad\quad V_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
+$V_{\mathrm{IPW}}^{\mathrm{meta^{*}}}$ $=\textstyle\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
+$V_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}$ $=\textstyle\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
 with within-site variance
-$\left\{\begin{array}[]{l}V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}\\ V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right],\end{array}\right$.
+$\left\{\begin{array}[]{l}\!V_{\mathrm{IPW}}^{(k)}\!=\!\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\!\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\!\tau_{k}^{2}\\ V_{\mathrm{AIPW}}^{(k)}\!=\!\mathbb{V}\big[\tau(X)\!\mid\!H=k\big]\!+\!\mathbb{E}\Big[\left(\frac{(Y\!-\!\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H\!=\!k\Big]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1\!-\!e_{k}(X)}\right)\!\mid\!H\!=\!k\right],\end{array}\right$.
 and $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$ the between-sites variance of the local ATEs.
@@
 In practice, treatment assignment frequently varies across sites due to differences in norms, resources, or clinical practices.
-To properly account for this heterogeneity, the global propensity score must be expressed as a weighted combination of the site-specific scores (see Appendix A.4).
-A first choice of weights are the density ratio weights (DW) $\omega_{k}^{\mathrm{DW}}$:
-$e(x)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(x)}{f(x)}}_{=\omega_{k}^{\mathrm{DW}}(x)}e_{k}(x),$ (3)
+This heterogeneity implies that the global propensity score takes the form of a weighted combination of the site-specific scores, that is $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$, see Appendix A.4.
+These weights $\{\omega_{k}\}_{k\in[K]}$ can be written as density ratio weights (DW):
+$e(x)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(x)}{f(x)}}_{=\omega_{k}}(x)e_{k}(x),$ (4)
 where $f_{k}$ and $f$ are the covariate densities locally and globally.
 Similar weights are used in transportability methods that reweight data to match a target population (Han et al., 2023, 2025; Guo et al., 2024), though here the objective is to recover the global propensity score across the super-population defined by the $K$ participating sites.
-Unfortunately, DW estimation in a federated setting requires modeling the $f_{k}$’s, which entails strong distributional assumptions and becomes challenging in high dimensions.
-Instead, we propose the Membership Weights (MW) $\omega_{k}^{\mathrm{MW}}$:
-$e(x)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X=x)}_{\displaystyle=\omega_{k}^{\mathrm{MW}}(x)}e_{k}(x),$ (4)
+Unfortunately, estimating DW $\omega_{k}$ in a federated setting requires modeling the $f_{k}$’s, which entails strong distributional assumptions and becomes challenging in high dimensions.
+Instead, we propose to rewrite the $\omega_{k}$’s as Membership Weights (MW) :
+$e(x)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X=x)}_{\displaystyle={\omega_{k}}(x)}e_{k}(x),$ (5)
 which represent the probability of site membership given the covariates.
@@
 We refer to Section 3.2 for more details on the federated estimation of MW and its advantages over DW.
-Equations 3 and 4 enable combining locally estimated propensity scores $e_{k}$ into a global propensity score using globally learned weights $\omega_{k}(x)$.
+Equations 4 and 5 enable combining locally estimated propensity scores $e_{k}$ into a global propensity score using globally learned weights $\omega_{k}(x)$.
 Building on this decomposition, we define our oracle Federated IPW and AIPW estimators (Fed-(A)IPW).
@@
 Under Assumptions 1, 2 and 4, we have:
-$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}],\quad\quad\quad\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}],$
+$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}],$
+$\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}],$
 with equality when the local propensity scores are identical across sites.
@@
 Second, our federated approach improves overlap compared with meta-analysis, as formalized below.
-Theorem 5 (Overlap improvement).
+Theorem 5.
+We compare the improvement in the overlap
 $0\leq\mathcal{O}_{\mathrm{global}}\leq\sum_{k=1}^{K}\rho_{k}\mathcal{O}_{k}.$
@@
 A key advantage of our approach is flexibility: sites can use different estimation methods tailored to local data or computational constraints.
-It also does not require Assumption 4: local scores may approach 0 or 1 provided the global score remains bounded away from these extremes.This, in particular, enables the integration of external control arms (FDA, 2023; EMA, 2023), where some sites have $e_{k}(X)=0$ for all control patients yet still contribute to the global analysis.
+It also does not require Assumption 4: local scores may approach 0 or 1 provided the global score remains bounded away from these extremes.This, in particular, enables the integration of external control arms (FDA, 2023b; EMA, 2023), where some sites have $e_{k}(X)=0$ for all control patients yet still contribute to the global analysis.
 Federated weights: density ratio vs.
 membership.
-Parametric density ratio weights $\omega_{k}^{\mathrm{DW}}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}$ can be implemented in a one-shot fashion by sharing local density parameters with the server, which then reconstructs the global mixture and computes the weights.
+Parametric density ratio weights $\hat{\omega}_{k}^{\mathrm{DW}}(x)=\rho_{k}\frac{\hat{f}_{k}(x)}{\hat{f}(x)}$ can be implemented in a one-shot fashion by sharing local density parameters with the server, which then reconstructs the global mixture and computes the weights.
 A common choice is to assume parametric covariate distributions (say, Gaussian), estimate $(\hat{\mu}_{k},\hat{\Sigma}_{k})$ locally, and transmit them once to the server.
@@
 Nonparametric density estimation would relax these assumptions but is statistically inefficient and does not yet have practical federated implementations.
-In contrast, our membership weights $\omega_{k}^{\mathrm{MW}}(x)=\mathbb{P}(H=k\mid X=x)$ can be learned with any probabilistic multiclass classifier trained federatively—for example, logistic regression for simplicity and interpretability, or neural networks to capture complex nonlinearities.
+In contrast, $\hat{\omega}_{k}^{\mathrm{MW}}(x)=\widehat{\mathbb{P}}(H\!=\!k\!\mid\!X\!=\!x)$, our membership weights, can be learned with any probabilistic multiclass classifier trained federatively—for example, logistic regression for simplicity and interpretability, or neural networks to capture complex nonlinearities.
 Such models are readily supported by modern FL algorithms and software libraries.
@@
 Remark (Missing values).
-Our approach can naturally accommodate missing data.
+Our approach can naturally handle missing data.
 Local propensity scores can be estimated consistently with logistic regression or random forests (Jiang et al., 2020; Josse et al., 2024).
@@
 Finally, doubly robust estimators with missing data can be obtained via non-parametric federated outcome models (Mayer et al., 2020).
+3.3 Accommodating Site Effects
+In practice, the site membership $H$ may affect potential outcomes through systematic differences (e.g., infrastructure, staff expertise, or measurement practices).
+When populations also differ across sites, ignoring such center effects can induce non-negligible confounding bias, since $H$ becomes associated with both $X$ and $Y$.
+In this section, we extend our framework to accommodate site effects when site-level covariates $Z$ (e.g., center-specific resources, equipment quality etc.) are observed and adequately capture these differences.
+We thus propose a relaxation of ˜2.
+Assumption 5 (Explained Site Effects).
+$\{Y(1),Y(0)\}\perp H\mid(X,Z).$
+Under ˜5, the effect of $H$ on $Y$ is fully mediated by $Z$, so that conditional on $(X,Z)$, there is no residual prognostic information carried by $H$.
+This assumption integrates naturally into our framework by incorporating $Z$ into the membership weights decomposition of Eq.
+(5):
+$e(X,Z)=\sum_{k=1}^{K}\mathbb{P}(H=k\mid X,Z)e_{k}(X).$
+In practice, we learn the resulting membership weights $\hat{\omega}_{k}^{\mathrm{MW}}(x,z)\!=\!\widehat{\mathbb{P}}(H\!=\!k\!\mid\!X\!=\!x,Z\!=\!z)$ as described in Section 3.2, with $Z$ included as additional covariates.
+Local propensity scores $e_{k}(X)$ remain functions of $X$ only because $Z$ is constant within each site.
+The AIPW outcome regressions are likewise extended to $\mu_{w}(X,Z)$.
 4 Experiments
-Synthetic data.
 (a) No local overlap
@@
 (d) Synthetic data: DGP A (top) and DGP B (bottom).
+4.1 Synthetic Data
 We consider $K=3$ sites and $d=10$ covariates.
@@
 1); meta-analysis IPW/AIPW with sample-size weighting (Meta-SW) (Def.
-2); and the one-shot inverse-variance weighted IPW estimator (1S-IVW) of Xiong et al.
+3); and the one-shot inverse-variance weighted IPW estimator (1S-IVW) of Xiong et al.
 (2023), evaluated under favorable conditions with shared propensity-score parameters across sites.
 For all estimators, propensity scores are fit via logistic regression and outcome models via linear regression, implying misspecification in the latter.
+Behavior under varying overlap.
 Figures 44 (top: DGP A; bottom: DGP B) summarize the results under the three overlap regimes.
@@
 In the Good local overlap setting (Figure 4), all methods are unbiased, but federated estimators achieve the smallest variance.
+Comparison to Xiong et al.
+(2023).
 Figure 3 considers a setting where all local propensity scores share a common subset of 5 of 10 logistic regression coefficients with data generated from DGP B.
@@
 Fed-MW remains unbiased and attains the lowest variance, matching that of 1S-IVW, even without access to the additional information about shared parameters.
+Handling site effets.
+We assess the benefit of adjusting for site effects when site-level covariates $Z$ are observed (Section˜3.3).
+We simulate $K=30$ sites, with heterogeneity arising both from diverse distributions of individual-level covariates $X$ (via covariate shift induced by $H\sim\mathrm{Logit}(\theta^{\top}X)$) and from site-level variation in outcomes captured by $Z$.
+Outcomes follow $Y=X^{\top}\beta+Z^{\top}\delta+\tau W+\varepsilon$, so that $H$ affects $Y$ only through $Z$.
+Federated methods incorporate $Z$ only in the membership weights, while meta-analysis adjusts only for $X$ as it naturally handles site effects.
+As shown in Table 1, incorporating $Z$ into our method eliminates confounding due to both covariate shift ($H\to X$) and mediated site effects ($H\to Z\to Y$).
+As a result, our Federated IPW/AIPW markedly improves MSE, especially for small $n_{k}$, where meta-analysis suffers from unstable local nuisance estimation.
 Additional simulations provided in Appendix E examine scenarios with more sites (Table 5), nonparametric estimation (Table 6), and local propensity model misspecification (Table 7), further confirming the robustness of our approach.
-Figure 2: Comparison to Xiong et al. (2023) (IPW).
+Estimator $n_{k}{=}20$ $n_{k}{=}60$ $n_{k}{=}100$
+IPW Oracle $(e^{\star}(X,Z))$ 0.853 0.443 0.091
+IPW Meta-SW $(\hat{e}(X))$ 5.176 0.777 0.590
+IPW Fed-MW (Logit, $\hat{e}(X,Z)$) 0.427 0.132 0.043
+IPW Fed-MW (Logit, $\hat{e}(X)$) 2.001 1.384 0.327
+IPW Fed-MW (NN, $\hat{e}(X,Z)$) 1.083 0.465 0.206
+IPW Fed-MW (NN, $\hat{e}(X)$) 2.100 1.418 0.356
+AIPW Oracle $(e^{\ast}(X,Z))$ 0.015 0.005 0.002
+IPW Meta-SW $(\hat{e}(X))$ 10.673 0.346 0.084
+AIPW Fed-MW (Logit, $\hat{e}(X,Z)$) 0.022 0.004 0.002
+AIPW Fed-MW (Logit, $\hat{e}(X)$) 2.026 1.348 0.323
+AIPW Fed-MW (NN, $\hat{e}(X,Z)$) 0.020 0.004 0.002
+AIPW Fed-MW (NN, $\hat{e}(X)$) 2.043 1.349 0.315
+Table 1: Mean Squared Error (MSE) of federated and meta-analysis estimators under Assumption 5.
+Figure 2: Comparison to Xiong et al. (2023) (IPW estimators).
 Figure 3: AIPW estimates on Traumabase.
-Real data.
-We analyze the multi-site Traumabase cohort (Mayer et al., 2020; Colnet et al., 2024) to estimate the effect of tranexamic acid on mortality across $K=14$ centers.
-The local datasets are highly imbalanced in site sizes (106 to 2,092 patients) and treatment arms (e.g., site 11: 4 treated vs.
+4.2 Real Data
+We analyze the multi-site Traumabase cohort (Mayer et al., 2020; Colnet et al., 2024) to estimate the effect of tranexamic acid on mortality across K=14K\!=\!14 centers.
+The local datasets are highly imbalanced in sizes (106 to 2,092 patients) and treatment arms (e.g., site 11: 4 treated vs.
 121 controls); there are 17 covariates and n=8,248n=8{,}248 patients in total, of whom 638 were treated.
 Covariates are standardized federatively by sharing site means and variances.
-We focus on AIPW estimators: local propensities $e_{k}$ are estimated via logistic regression; outcome models $\mu_{1},\mu_{0}$ are trained with FedAvg logistic regression (5,000 rounds, 1 local epoch, step size $\eta=0.1$); MW are learned with a FedAvg neural network (one hidden layer, 128 units), while DW are based on Gaussian density estimation.
-Competitors include Meta-SW (same local models, computed only on sites with enough treated units—where the number of treated observations exceeds the covariate dimension) and a centralized AIPW benchmark obtained using the R package grf’s probability_forest function (Tibshirani et al., 2018) on the pooled data.
+We focus on AIPW estimators: local propensities $e_{k}$ are estimated using the R package grf’s probability_forest function (Tibshirani et al., 2018) on large sites, and logistic regressions on smaller ones; outcome models $\mu_{1},\mu_{0}$ are trained with FedAvg logistic regression (5,000 rounds, 1 local epoch, step size $\eta=0.1$); MW are learned with a FedAvg neural network (NN, one hidden layer, 128 units) and a federated logistic regression.
+We also report a centralized random forest (RF) MW as a strong benchmark.
+DW are based on Gaussian density estimation.
+Competitors include Meta-SW (same local models, computed only on sites with enough treated units—where the number of treated observations exceeds the covariate dimension) and a centralized AIPW benchmark obtained using probability_forest function on the pooled data.
 Empirical confidence intervals are constructed from 150 bootstrap resamples.
-Figure 3 shows that our federated estimators closely match the centralized benchmark and exhibit lower variance than Meta-SW, which in this application departs from the centralized estimate.
-Among federated methods, Fed-MW is closest to the nonparametric centralized estimator—reflecting its more flexible modeling of propensities and membership—whereas Fed-DW is less stable, likely due to noisy per-site covariance estimates for a 17-dimensional Gaussian ($\approx 17^{2}$ parameters) in small sites.
+Figure 3 shows that our federated MW estimators closely match the centralized benchmark and exhibit lower variance than Meta-SW, which in this application departs from the centralized estimate.
+Among federated methods, Fed-MW are closest to the nonparametric centralized estimators—reflecting their more flexible modeling of propensities and membership—whereas Fed-DW is less stable, likely due to noisy per-site covariance estimates for a 17-dimensional Gaussian ($\approx 17^{2}$ parameters) in small sites.
 5 Conclusion, Limitations and Future work
-We propose a theoretically grounded framework for federated causal inference that leverages membership weights to construct valid pseudo-populations across silos.
-These weights, estimated via flexible parametric or nonparametric models, improve overlap and yield more stable ATE estimates, without sharing raw data.
-Our framework accommodates heterogeneous local propensity score estimation strategies, supports external control arms, and remains robust to even extreme local treatment–control imbalances.
-Although sufficiently large per-site datasets are still needed—particularly in high-dimensional settings—our approach is especially well suited to a moderate number of large silos, where FL most effectively increases effective sample size.
-Promising avenues for future work include principled handling of site effects (i.e., relaxing Assumption 2) and extending our framework to CATE estimation.
+We propose a theoretically grounded framework for federated causal inference that leverages membership weights to construct valid causal estimates across silos.
+These weights, estimated via flexible parametric or nonparametric models, improve overlap and yield more stable ATE estimates.
+Our framework accomodates arbitrary heterogeneity in local propensity score functions, does not require local overlap to hold, enabling the inclusion of small sites and external control arms, without sharing individual-level data.
+Our approach can also be naturally extended to handle some form of site effects.
+Although sufficiently large per-site datasets are still needed—particularly in high-dimensional settings—to learn local propensity scores accurately, our approach is especially well suited to a moderate number of large silos, where federated learning most effectively increases effective sample size.
+A promising avenue for future work is to extend our framework to CATE estimation.
 While the ATE remains central in econometrics, biomedicine, and public policy, considering CATE is important for moving towards personalization.
@@
 Proof.
-Under Assumptions 1 and 2, the variables in $X_{i}$ are a sufficient adjustment set, so that the ATE is identifiable as
+Under Assumptions 1 and 2, the variables in $X$ form a sufficient adjustment set. We identify the ATE as follows:
 $\tau$ $=\mathbb{E}\left[Y(1)-Y(0)\right]$
-$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X\right]-\mathbb{E}\left[Y(0)\mid X\right]\right]$
-$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,W=1\right]-\mathbb{E}\left[Y(0)\mid X,W=0\right]\right]$ Assumptions 1 and 2
-$=\mathbb{E}\left[\mathbb{E}\left[Y\mid X,W=1\right]-\mathbb{E}\left[Y\mid X,W=0\right]\right]$ SUTVA
+$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,H\right]-\mathbb{E}\left[Y(0)\mid X,H\right]\right]$ Law of Iterated Expectations
+$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,H,W=1\right]-\mathbb{E}\left[Y(0)\mid X,H,W=0\right]\right]$ Assumption 1
+$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,W=1\right]-\mathbb{E}\left[Y(0)\mid X,W=0\right]\right]$ Assumption 2 and Lemma 4.3 Dawid [1979]
+$=\mathbb{E}\left[\mathbb{E}\left[Y\mid X,W=1\right]-\mathbb{E}\left[Y\mid X,W=0\right]\right]$ Consistency
 $=\mathbb{E}\left[\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\right]$ Assumption 3
@@
 Proof.
-Unbiasedness. We have:
-$\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right]$
-$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\cdot\mathbb{1}_{[H_{i}=k]}\mid n_{1},\dots,n_{K}\right]\right]$
-$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[n_{k}\right]\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$
-$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$ (6)
-Let us focus on the first term of the sum:
-$\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]$ $=\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid H_{i}=k\right]$ SUTVA
-$=\mathbb{E}\left[\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid X_{i},H_{i}=k\right]\right]$
-$=\mathbb{E}\left[\frac{\mathbb{E}\left[W_{i}\mid X_{i},H_{i}=k\right]\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$ Assumptions 1 and 2
-$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$
-$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{\mathbb{E}\left[e_{k}(X_{i})\mid X_{i},H_{i}=k\right]}\right]$ Definition of $e(X_{i})$
-$=\mathbb{E}\left[\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]\mathbb{E}\left[\frac{e_{k}(X_{i})}{\mathbb{E}\left[e_{k}(X_{i})\right]}\mid X_{i}\right]\right]$
-$=\mathbb{E}\left[Y_{i}(1)\mid H_{i}=k\right]$
-Similarly, we have $\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]=\mathbb{E}\left[Y_{i}(0)\mid H_{i}=k\right]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
-For the AIPW estimator, following the same steps as in section˜A.2, we have
-$\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
-$=\sum_{k=1}^{K}\rho_{k}\Bigg(\mathbb{E}\left[\mu_{1}(X_{i})-\mu_{0}(X_{i})\mid H_{i}=k\right]$
-$\quad\quad+\mathbb{E}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\mid H_{i}=k\right]\Bigg)$
-$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\tau(X_{i})\mid H_{i}=k\right]+0+0\right)$
-$=\tau$
+Unbiasedness. We show the unbiasedness of the first term of the IPW estimator. The derivation for the second term is symmetric.
+$\mathbb{E}\left[\frac{WY}{e(X)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\frac{WY(1)}{e(X)}\middle|X,H\right]\right]$ SUTVA
+$=\mathbb{E}\left[\frac{\mathbb{E}\left[W\mid X,H\right]\mathbb{E}\left[Y(1)\mid X,H\right]}{e(X)}\right]$ Assumption 1
+$=\mathbb{E}\left[\frac{e_{H}(X)\mathbb{E}\left[Y(1)\mid X,H\right]}{e(X)}\right]$ $\text{Definition of local score }e_{H}(X)$
+$=\mathbb{E}\left[\frac{e_{H}(X)\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}\right]$ Assumption 2
+$=\mathbb{E}_{X}\left[\frac{\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}\mathbb{E}_{H}\left[e_{H}(X)\mid X\right]\right]$ $\text{Iterated Expectation over }H$
+$=\mathbb{E}_{X}\left[\frac{\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}\sum_{k=1}^{K}\mathbb{P}(H=k\mid X)e_{k}(X)\right]$
+$=\mathbb{E}_{X}\left[\frac{\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}e(X)\right]$ $\text{Definition of global }e(X)$
+$=\mathbb{E}\left[\mathbb{E}[Y(1)\mid X]\right]=\mathbb{E}[Y(1)]$
+Note that the step applying Assumption 2 relies on the consequence that $\mathbb{E}[Y(1)\mid X,H]=\mathbb{E}[Y(1)\mid X]$ (which is derived from Assumption 2 combined with Assumption 1).
+Similarly, we have $\mathbb{E}\left[\frac{(1-W)Y}{1-e(X)}\right]=\mathbb{E}[Y(0)]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\tau$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
+For the AIPW estimator, we note that $\mathbb{E}[W(Y-\mu_{1}(X))/e(X)]=0$ following the same logic as above (replacing $Y(1)$ with the residual $Y(1)-\mu_{1}(X)$ which has mean zero given $X$). Thus:
+$\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\mu_{1}(X)-\mu_{0}(X)+\frac{W(Y-\mu_{1}(X))}{e(X)}-\frac{(1-W)(Y-\mu_{0}(X))}{1-e(X)}\right]$
+$=\mathbb{E}[\mu_{1}(X)-\mu_{0}(X)]+0-0$
+$=\mathbb{E}[\tau(X)]=\tau$
 Variance.
@@
 Algorithm 1 summarizes the generic Federated Averaging procedure for training any differentiable parametric model for membership-weight estimation.
+0:$K$ sites, $E$ local steps, $\eta$ learning rate, $T$ communication rounds, $B$ batch size
+1: Initialize global parameters $\Theta_{0}$
+2:for $t=1$ to $T$ do
+3:  for each client $k\in[1,\dots,K]$ in parallel do
+4:   $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}$)
+5:  end for
+6:  $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t+1}^{(k)}$ // FedAvg aggregation
+7:end for
+8:Function LocalUpdate($k$, $\Theta$):
+9:for $e=1$ to $E$ do
+10:  Sample mini-batch $\mathcal{B}_{k}$ of size $B$ from local data $\mathcal{D}_{k}$
+11:  Compute predictions $\omega_{k}(X_{i};\Theta)$ for $i\in\mathcal{B}_{k}$
+12:  Compute gradient $\nabla\ell(\Theta;\mathcal{B}_{k})$ of cross-entropy loss
+13:  $\Theta\leftarrow\Theta-\eta\nabla\ell(\Theta;\mathcal{B}_{k})$
+14:end for
+15:Return $\Theta$
 Algorithm 1 Federated Learning of Membership Weights with a Parametric Model
-1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ communication rounds, $B$ batch size
-2:Initialize global parameters $\Theta_{0}$
-3:for $t=1$ to $T$ do
-4:  for each client $k\in[1,\dots,K]$ in parallel do
-5:   $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}$)
-6:  end for
-7:  $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t+1}^{(k)}$ // FedAvg aggregation
-8:end for
-9:LocalUpdate($k$, $\Theta$):
-10:for $e=1$ to $E$ do
-11:  Sample mini-batch $\mathcal{B}_{k}$ of size $B$ from local data $\mathcal{D}_{k}$
-12:  Compute predictions $\omega_{k}(X_{i};\Theta)$ for $i\in\mathcal{B}_{k}$
-13:  Compute gradient $\nabla\ell(\Theta;\mathcal{B}_{k})$ of cross-entropy loss
-14:  $\Theta\leftarrow\Theta-\eta\nabla\ell(\Theta;\mathcal{B}_{k})$
-15:end for
-16:return $\Theta$
 With a suitable choice of learning rate $\eta$ and a moderate number of local steps $E$ per round, Algorithm 1 converges to the same solution as centralized training as the number of communication rounds $T\to\infty$ [Stich, 2019, Khaled et al., 2020, Li et al., 2019].
@@
 $\hat{\omega}_{k}(X_{i})=\omega_{k}(X_{i};\widehat{\Theta}^{\mathrm{fed}}),\qquad k=1,\dots,K.$
+Note that this federated procedure remains valid for learning $H|X$ despite each site $k$ locally holding only one label ($H=k$).
 B.3 Example: Multinomial Logistic Regression
@@
 Appendix C Simulation Details
-The parameters common to all settings are shown in Table 1, where $\gamma_{2}^{\mathrm{(weak)}}=[-2.5,-1,-0.15,-0.15,0,-0.15,-1,-0.15,-0.15,0]$ and $\gamma_{2}^{\mathrm{(good)}}=[-.05,-.1,-.05,-.1,.05,-.1,-.05,-.1,.05,-.1]$.
+The parameters common to all settings are shown in Table 2, where
+$\gamma_{2}^{\mathrm{(weak)}}=[-2.5,-1,-0.15,-0.15,0,-0.15,-1,-0.15,-0.15,0]$
+and
+$\gamma_{2}^{\mathrm{(good)}}=[-.05,-.1,-.05,-.1,.05,-.1,-.05,-.1,.05,-.1]$
 Parameter Center 1 Center 2 Center 3
@@
 $e_{k}$ $\mathrm{Logistic}(x,\gamma_{k})$
-$\gamma_{k}$ $\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$
-No overlap: not logistic
-(only controls)
-Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$
-Good overlap: $\gamma_{2}^{\mathrm{(good)}}$
-$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
+$\gamma_{k}$ $\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$ No overlap: not logistic (only controls) Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$ Good overlap: $\gamma_{2}^{\mathrm{(good)}}$ $\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
 Table 2: Common simulation parameters.
@@
 Table 4: Simulation parameters specific to DGP B.
+Real data.
+We classified a site as ‘large’ if it had at least $5\times d$ observations in each treatment arm (minimum 85 per arm in our setting).
+For such sites, we estimated $e_{k}$ with probability forests with $2,000$ trees; otherwise, we used logistic regression.
+We estimated Fed-MW RF with a centralized random forest for $P(H|X)$ with $500$ trees.
 Appendix D Overlap Improvement
@@
 As expected, the AIPW estimator retains its double-robustness property even under misspecification of the outcome model, leading to consistently low bias and MSE across settings.
+Appendix F Additional Details on Handling Site Effects
+F.1 Federated Tests for Site Effects
+Assumption 2 posits that site membership $H$ provides no information about potential outcomes conditional on covariates $X$ and treatment $W$.
+This assumption is testable by fitting a federated regression model of $Y$ against $(X,H)$, and testing whether the coefficients for $H$ are jointly zero.
+In federated learning, this test is implemented via one-shot sharing of sufficient statistics: each site transmits its local covariance matrix $\hat{\Sigma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}X_{i}^{\top}X_{i}$ and vector $\hat{\gamma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}X_{i}^{\top}Y_{i}$.
+The server aggregates these to compute a joint F-test or Wald test on the $H$ block, yielding a $p$-value.
+We recommend adapting the estimation strategy based on this test’s result.
+WWXXYYHHZZ
+Figure 6: Graphical model for multi-site observational data with site covariates $Z$ mediating the site effect.
+F.2 Experimental Details
+We provide the simulation parameters used in LABEL:table:site-effects.
+We simulate $K=30$ sites, each with $n_{k}\in\{20,60,90\}$ observations.
+Covariates $X\in\mathbb{R}^{3}\sim\mathcal{N}(0,I_{3})$ are drawn i.i.d.
+and allocated to sites via $H\sim\text{Logit}(\theta^{\top}X)$ to induce covariate shift.
+Site-level covariates $Z\in\mathbb{R}^{3}$ (site effects) are drawn from site-specific Gaussians: for each site $k$, a centroid $\mu_{k}\sim\mathcal{N}([1,1,1],I_{3})$ is drawn, then $Z\mid H=k\sim\mathcal{N}(\mu_{k},I_{3})$.
+Treatment is assigned by site-specific logistic models.
+Outcomes are linear: $Y=X^{\top}\beta+Z^{\top}\delta+\tau W+\varepsilon$, with $\varepsilon\sim\mathcal{N}(0,1)$, $\tau=10$, and $\beta=\delta=[-2,-1.33,-0.66]$.
+We estimate membership probabilities $P(H\mid\cdot)$ using multinomial logistic regression or a two-layer neural network (NN), trained on either $(X)$ or $(X,Z)$.
+Local propensities $\hat{e}_{k}(X)$ are fitted via logistic regression, and federated outcome models $\hat{\mu}_{w}(X,Z)$ via OLS.
+Membership weights are aggregated to form $\hat{e}(x,z)=\sum_{k}\hat{P}(H{=}k\mid x,z)\hat{e}_{k}(x)$.
+IPW and AIPW estimators are computed as $\hat{\tau}_{\mathrm{IPW}}=\frac{1}{n}\sum_{i}\frac{W_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{1}{n}\sum_{i}\frac{(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}$ and $\hat{\tau}_{\mathrm{AIPW}}=\frac{1}{n}\sum_{i}\big[\frac{W_{i}(Y_{i}-\hat{\mu}_{1})}{\hat{e}(X_{i})}-\frac{(1-W_{i})(Y_{i}-\hat{\mu}_{0})}{1-\hat{e}(X_{i})}+(\hat{\mu}_{1}-\hat{\mu}_{0})\big]$.
+Meta-analysis estimators adjust only for $X$, as averaging over $H$ inherently stratifies by site.
+F.3 On Unexplained Site Effects
+If site-level covariates $Z$ are unavailable or insufficient to explain all site effects (e.g., due to a direct $H\to Y$ edge, see Figure˜7), the core identification assumption remains Assumption 1: $\left\{Y(1),Y(0)\right\}\perp W\mid(X,H)$).
+In this regime, nuisance functions—propensity scores $e(X,Z,H)$ and outcome models $\mu_{w}(X,Z,H)$—must be learned federatively using parametric models that explicitly include site fixed effects ($H$).
+Including $Z$ in these models helps to reduce the variance of the nuisance functions.
+Note that the theoretical and practical advantage of federated learning over meta-analysis is diminished in this setting, as meta estimation effectively stratifies inference on $H$.
+WWXXYYHHZZ
+Figure 7: Graphical model for multi-site observational data where site effects are not fully explained by $Z$ (direct $H\to Y$ edge).